### v1

# Expert-elicitation method for non-parametric joint priors using normalizing flows

Abstract
We propose an expert-elicitation method for learning non-parametric joint prior distributions using normalizing flows. Normalizing flows are a class of generative models that enable exact, single-step density evaluation and can capture complex density functions through specialized deep neural networks. Building on our previously introduced simulation-based framework, we adapt and extend the methodology to accommodate non-parametric joint priors. Our framework thus supports the development of elicitation methods for learning both parametric and non-parametric priors, as well as independent or joint priors for model parameters. To evaluate the performance of the proposed method, we perform four simulation studies and present an evaluation pipeline that incorporates diagnostics and additional evaluation tools to support decision-making at each stage of the elicitation process.
keywords:
prior elicitation, expert knowledge, joint prior distribution, normalizing flows, non-parametric priors
1 Introduction
The Bayesian paradigm offers the possibility to incorporate prior knowledge into a statistical model through the specification of prior distributions. This possibility is a central advantage of the Bayesian paradigm [50], yet it also presents one of its most challenging aspects [68, 65, 70]. In the following, we define prior knowledge as the expertise provided by a domain expert — an individual with extensive knowledge of a specific subject matter [23]. This knowledge can be represented in various forms, but to integrate it into a Bayesian model, we need to translate it into a formal mathematical language that can be expressed as a prior distribution over the model parameters [59, 57, 48, 28].
A whole field of research, commonly referred to as (expert) prior elicitation, has emerged around the question of how to gather expert knowledge and translate it into appropriate prior distributions [69, 50, 23]. Garthwaite et al [28] identified four key stages in a prior elicitation process:
  1. 1.
    Setup stage: In this stage, the problem is defined, an expert is selected, and the quantities to be elicited from the expert (referred to as target quantities) are determined;
  2. 2.
    Elicitation stage: Here, the target quantities are queried from the expert using specific elicitation techniques, resulting in what we call elicited statistics;
  3. 3.
    Fitting stage: This involves fitting a (potentially joint) probability distribution based on the expert-elicited statistics;
  4. 4.
    Evaluation stage: Finally, the adequacy of the fitted probability distribution is assessed in collaboration with the expert.
In this context, elicitation methods aim to provide a systematic and formal procedure for deriving prior distributions based on expert-elicited statistics. Early elicitation methods primarily tackled the problem by seeking analytical solutions, such as conjugate models or problem-specific transformation functions resulting in highly model-specific prior elicitation methods [see 50, for a recent review]. This model dependence is partly due to the use of direct elicitation techniques, which involve asking experts directly about model parameters [69, 23]. This approach is problematic not only due to its inherent model dependence but also because model parameters are often challenging for domain experts to interpret meaningfully. Consequently, the quality of expert input can be questionable. In response to these limitations, substantial efforts have been made to develop methods that enable more flexible and interpretable prior specification for domain experts.
[…]
The methods introduced so far that use this simulation-based approach focus on learning independent, parametric priors $p(\theta\mid\lambda)$, parameterized by a set of prior hyperparameters $\lambda$ [38, 67, 47]. While assuming independent priors for model parameters can be reasonable in some cases due to theoretical considerations or model constraints, this approach may not be sufficient for more complex or high-dimensional problems [30, 32, 68]. In such cases, it becomes important to account for the joint distribution of model parameters to capture dependencies and interactions that reflect the underlying structure more accurately. Similarly, the use of parametric prior distribution families can be well-justified in some cases, but may lead to model misspecification in others, prompting the need for non-parametric approaches. Elicitation methods that focus on non-parametric priors are less studied, but examples include the use of Gaussian processes [54] and quantile-parameterized distributions [59, 60].
Building in particular on the work of Hartmann et al [38], da Silva et al [67], Manderson and Goudie [47], we previously introduced a simulation-based framework with a highly modular structure that closely aligns with the elicitation process described by Garthwaite et al [28]. The advantage of this modular structure is its flexibility, allowing for easy adaptation to different types of methods. In our previous work [6], we demonstrated how to use this framework to learn parametric prior distributions for the model parameters. In this paper, we show how the same simulation-based framework can be used to learn flexible (i.e., non-parametric) joint priors for the model parameters with only minor adjustments to the workflow. In both approaches, we employ mini-batch stochastic gradient descent as the optimization method. However, our framework supports different optimization methods thanks to its modular design.
Main Contributions
Our motivation is two-fold: First, we aim to promote the development of a unifying, highly modular framework that supports a wide range of prior elicitation methods. We believe that such a framework would benefit both users and developers of prior elicitation methods by providing structure and clarity in a field currently populated with numerous coexisting methods, which can make comparisons challenging. Second, we want to demonstrate how our simulation-based framework, introduced in Bockting et al [6], can serve as a flexible approach for learning either parametric or non-parametric, as well as independent or joint, priors with only minor adjustments. While our previous work focused on learning parametric, independent priors, the present work emphasizes learning non-parametric, joint priors.
2 Methodology
In this paper, we propose an elicitation method that builds on and extends our recently introduced simulation-based framework [6]. The overall idea of this framework closely resembles the approach of prior predictive checks [32]: We simulate from the joint model $p(\theta,y)$ and assess how well the resulting prior predictions align with the expert’s expectations. If there is a discrepancy between the expert’s expectations and the simulations, the prior specification needs to be adjusted accordingly. Figure 1 provides a graphical representation of the framework and Table 1 offers a summary of the symbols and notation used in the following sections. The general workflow of our framework can be summarized as follows:
  1. 1.
    Define the generative model: Define the generative model including dimensionality and parameterization of prior distribution(s). (Setup stage; Section 2.1)
  2. 2.
    Identify variables and elicitation techniques for querying expert knowledge: Select the set of variables to be elicited from the domain expert (target quantities) and determine which elicitation techniques to use for querying the selected variables from the expert (elicited statistics). (Setup stage; Section 2.2)
  3. 3.
    Elicit statistics from expert and simulate corresponding predictions from the generative model: Sample from the generative model and perform all necessary computational steps to generate model predictions (model-elicited statistics) corresponding to the set of expert-elicited statistics. (Elicitation stage; Section 2.3)
  4. 4.
    Evaluate consistency between expert knowledge and model predictions: Evaluate the discrepancy between the model- and expert-elicited statistics via a multi-objective loss function. (Fitting stage; Section 2.4)
  5. 5.
    Adjust prior to align model predictions more closely with expert knowledge: Use mini-batch stochastic gradient descent to adjust the prior so as to reduce the loss. (Fitting stage; Section 2.4)
  6. 6.
    Find prior that minimizes the discrepancy between expert knowledge and model predictions: Repeat steps 2 to 5 iteratively until a prior is found that minimizes the discrepancy between the model and expert-elicited statistics. (Fitting stage; Section 2.4)
  7. 7.
    Evaluate the learned prior distributions: Run the learning algorithm (steps 2 to 6) multiple times to obtain a set of prior distributions that can equally well represent the expert data. Select a plausible prior distribution in consultation with the domain expert or apply model averaging techniques. (Evaluation stage; Section 2.5)
The following sections discuss each step in greater detail.
Figure 1: Graphical illustration of our simulation-based prior-elicitation framework. The process begins by identifying target quantities to be elicited from the domain expert and selecting appropriate elicitation techniques, which result in (expert-)elicited statistics. Next, predictions are simulated from the generative model by sampling from the prior $p_{\lambda}(\theta)$ and computing the corresponding model-implied target quantities and elicited statistics. The consistency between model-and expert-elicited statistics is assessed using a loss function $L_{m}$, where $\alpha_{m}$ is the weight of the $m^{\text{th}}$ loss component. Hyperparameters $\lambda$ that define the learned prior are adjusted based on this evaluation to reduce the loss and align model predictions more closely with expert knowledge. This iterative process continues until a prior is found that minimizes the discrepancy between model and expert-elicited statistics. In the updating rule, $\delta$ refers to the step size.
Symbol Description
Simulation-based prior elicitation
$y_{n}$ data $n=1,\ldots,N$
$\theta_{k}$ model parameter $k=1,\ldots,K$
$\lambda$ model hyperparameter
$p(y\mid\theta)$ likelihood
$p(\theta\mid\lambda)$ prior parameterized by $\lambda$
$p(y)$ prior predictive distribution
$z_{i}=c_{i}(\lambda)$ simulated model target quantity defined by function $c_{i}$
$\hat{z}_{i}$ expert representation of the target quantity
$t_{m}=f_{j}(z_{i})$ model-elicited statistic defined by elicitation technique $f_{j}$
$\hat{t}_{m}$ expert-elicited statistic
$L_{m}(t_{m}(\lambda),\hat{t}_{m})$ loss component
$L(\lambda)=\sum_{m=1}^{M}\alpha_{m}L_{m}(t_{m}(\lambda),\hat{t}_{m})$ total loss as weighted sum of $L_{m}$ with weights $\alpha_{m}$
$p_{\lambda_{1},\ldots,\lambda_{R}}(\theta)=\sum_{r=1}^{R}w_{r}\cdot p_{\lambda_{r}}(\theta)$ averaged prior across $R$ replications and weighted by $w_{r}$ with $r=1,\ldots,R$
$\Delta_{r}(L)=L(\lambda_{r})-L(\lambda_{\min})$ difference between total loss of replication $r$ and replication with minimum total loss
$w_{r}=\frac{\exp\{-\gamma\Delta_{r}(L)\}}{\sum_{v=1}^{V}\exp\{-\gamma\Delta_{v}(L)\}}$ weights used for model averaging with $\gamma$ being a scaling factor; in all simulation studies $\gamma=1.$
Normalizing flow
$u\sim p_{U}(u)$ samples from base distribution
$g_{\lambda}(\cdot)$ generator function
$u=g_{\lambda}(\theta)$ normalizing direction
$\theta=g^{-1}_{\lambda}(u)$ generative direction
$g_{\lambda}=g_{\lambda_{H}}\circ\cdots\circ g_{\lambda_{1}}$ composition of $H$ coupling layers
Table 1: Notation and symbols used in this paper.
2.1 Prior Specification & Learning
In our previous work [6], we introduced the simulation-based framework along with a method for learning a set of independent parametric prior distributions, $p(\theta_{k}\mid\lambda_{k})$, for the model parameters $\theta_{k}$ with $k=1,\ldots,K$. In this paper, we extend our framework by introducing a method to learn a joint non-parametric prior distribution $p_{\lambda}(\theta)=p_{\lambda}(\theta_{1},\ldots,\theta_{K})$ over all model parameters $\theta$. In both the parametric and non-parametric approaches, the objective of the optimization process is to learn the hyperparameters $\lambda$ that minimizes the discrepancy between the model simulations and the expert expectations. However, the interpretation of these hyperparameters differs between approaches, which we emphasize by using a different notation for the prior distributions. In the non-parameteric approach, we employ normalizing flows (NFs) to induce a flexible family of prior distributions which entail specialized deep neural networks with trainable parameters $\lambda$ [44]. Thus, in the non-parametric approach, $\lambda$ represents the weights of the deep neural networks within the NFs, whereas in the parametric approach, $\lambda$ denotes the parameters of the associated prior distribution families (i.e., the prior hyperparameters).
Normalizing Flows
To learn a non-parametric joint prior for the model parameters, we use normalizing flows (NFs), a family of deep generative models that offers tractable density evaluation [44]. NFs transform a simple probability distribution, the base distribution $p(u)$, into a complex target distribution — in our case, the joint prior $p_{\lambda}(\theta)$ — through a series of invertible and differentiable mappings $g_{\lambda}$, parameterized by $\lambda$ [44]. The composition of invertible functions $g_{\lambda}$ yields an explicit form of the density $p_{\lambda}(\theta)$ via the change of variables formula
$p_{\lambda}(\theta)=p(u=g_{\lambda}(\theta))\mid\det g_{\lambda}^{\prime}(\theta)\mid,$
where $g_{\lambda}^{\prime}(\theta)=\frac{\partial}{\partial\theta}g_{\lambda}(\theta)$ is the Jacobian matrix of $g_{\lambda}$ at $\theta$ and $\mid\det g_{\lambda}^{\prime}(\theta)\mid$ is its absolute determinant. This formula allows us to easily evaluate the prior density at arbitrary points of interest $\theta$ [20]. Obtaining samples from the joint prior can be achieved by first sampling from the base distribution, $p(u)$, and then applying the inverse $g^{-1}_{\lambda}(u)$ to obtain samples from $p_{\lambda}(\theta)$:
$\theta=g^{-1}_{\lambda}(u)\sim p_{\lambda}(\theta)\quad\text{for}\quad u\sim p(u).$
Typically, $p(u)$ is chosen to be a simple density, such as a standard Gaussian or a uniform distribution [44]. Figure 2 depicts the use of NFs for learning a non-parameteric prior distribution within our simulation-based framework.
Figure 2: Learning non-parametric prior distributions using normalizing flows within our simulation-based framework. Normalizing flows transform a simple probability distribution, the base distribution $p(u)$, into a complex target distribution — in our case, the joint prior $p_{\lambda}(\theta)$ — through a series of invertible and differentiable mappings $g_{\lambda}$, parameterized by $\lambda$.
For NFs to be practical, the mapping $g_{\lambda}$ should be easily invertible, computationally efficient, and sufficiently expressive to model any target distribution of interest [44]. Several types of NFs that meet these requirements have been proposed, including affine [19] and spline coupling flows [22]. A detailed discussion and introduction to coupling flows can be found in various sources [21, 19, 20, 62, 44, 58, 7]. Due to their simple architecture, affine coupling flows are easy to invert and computationally efficient, although they have restricted expressiveness [7]. However, the expressiveness of a coupling flow can be increased by stacking multiple affine coupling layers, $g_{\lambda}=g_{\lambda_{H}}\circ\cdots\circ g_{\lambda_{1}}$ with $h=1,\ldots,H$ coupling blocks [20]. In fact, Draxler et al [21] prove that affine coupling flows are distributional universal approximators, capable of representing any target distribution despite their seemingly restrictive architecture. In line with these considerations and the strong empirical performance of affine coupling flows, we focused on this type of normalizing flows in our simulation studies (see Section 4), using the implementation provided by BayesFlow [63].
Finally, it is worth noting that, since our focus is solely on the mapping from the base distribution to the target distribution (i.e., the generative direction of NFs), virtually any other generative model family could be used to learn how to generate samples from a joint prior. Examples of common generative architecture alternatives include diffusion models [9] and flow matching [45]. However, we used NFs here because they perform well in practice and are computationally efficient. In particular, they do not require solving differential equations during sampling [21].
2.2 Target Quantities & Elicitation Techniques
A key task in the setup stage of the elicitation process is the selection of target quantities and elicitation techniques to effectively gather the expert information. Two main criteria should guide this selection: interpretability and informativeness [50, 14, 67, 28].
Selecting Target Quantities
Interpretability refers to the selection of target quantities that allows domain experts to meaningfully express their knowledge based on their experience and expertise [29]. Target quantities on the same scale as the outcome variable (cf., observable quantities) are considered to be highly interpretable [67, 47]. In contrast, target quantities that refer to model parameters are more challenging to interpret, especially when the outcome variable is not on the same scale as the model parameters [40, 41]. Therefore, it is sometimes advocated asking experts exclusively about observable quantities [40]. We share the view of Denham and Mengersen [15] that the appropriateness of either approach depends on the specific problem being addressed and the type of expert providing the information.
[…]
Selecting Elicitation Techniques
Once the set of target quantities has been selected, we need to determine how to elicit this information from the expert, which requires choosing an appropriate elicitation technique. While experts can be asked to provide the full distribution of a target quantity (e.g., in the form of a histogram [39]), another common technique is to focus on specific summary statistics [52]. Research on prior elicitation suggests that experts can reasonably be asked to describe a target quantity using proportions, the mode, the median, and generally quantiles [28, 69, 56], with quantile-based elicitation techniques being particularly recommended [40]. In contrast, asking about the mean is less advisable when the distribution is skewed [61], and similarly, querying variances is in general discouraged [28, 40]. Furthermore, assessing dependencies between variables is challenging, making the elicitation of joint priors especially difficult [28]. Elicitation of correlations in the context of joint priors has been studied among others by Gokhale and Press [33], Clemen et al [11], and Dickey et al [17] as reviewed by Mikkola et al [50].
In our elicitation method, we represent an elicitation technique as a function $f_{j}$ of a target quantity $z_{i}$. The resulting set of elicited target quantities is referred to as elicited statistics,  $\{t_{m}(\lambda)\}$, where $t_{m}(\lambda)=f_{j}(z_{i})$ and $m$ refers to the corresponding $i\times j$ combination. In the following, we sometimes keep the dependency on $\lambda$ implicit and simply write $t_{m}$ instead of $t_{m}(\lambda)$.
Sensitivity Analysis
In addition to interpretability, the second criterion for selecting target quantities and elicitation techniques is informativeness, which refers to the relevance of the elicited statistics for learning the prior distributions. Evaluating this criterion is nontrivial because the relationship between the elicited statistics and the model (hyper-)parameters is complex and often analytically intractable. One computational approach to asses informativeness is to conduct a sensitivity analysis [16]. In this approach, we systematically vary one aspect of the prior distribution at a time while keeping all other aspects constant. For each specific change in the prior, we run the generative model in forward mode by sampling from the prior, then from the likelihood, and subsequently computing the target quantities and elicited statistics. Sensitivity analysis enables us to assess how changes in one aspect of the prior distribution impact each elicited statistic. If we observe no variation in a particular elicited statistic for a given prior variation, this statistic does not provide any useful information for the learning algorithm to determine the corresponding aspect of the prior. This approach can help inform the selection of elicited statistics and identify which aspects of the prior are most likely to remain unidentifiable given the current set of expert information.
2.3 Model-implied & Expert-elicited Statistics
After determining the set of target quantities and corresponding elicitation techniques, we can proceed to gather information from the expert through specific prior elicitation protocols [35, 12, 2]. This process may be preceded by a training phase to familiarize the expert with the elicitation tasks [69]. In prior elicitation research, an entire subfield has focused on determining how to conduct optimal expert interviews, taking into account factors such as cognitive biases and heuristics. While it is beyond the scope of this paper to review this literature, several comprehensive reviews are available [e.g., 50, 69, 23]. At this point, we will skip the actual interrogation process and assume that the expert information has been successfully gathered, providing us with a set of expert-elicited statistics, denoted by $\hat{t}_{m}$.
Given the set of expert-elicited statistics, $\{\hat{t}_{m}\}$, we can evaluate the discrepancy between the expert expectations and the model-implied statistics, $\{t_{m}\}$. The latter are computed by simulating from the model. This proceeds as follows: We start with an arbitrary initial prior distribution from which we sample model parameters $\theta^{(s)}\sim p_{\lambda_{0}}(\theta)$, where $\lambda_{0}$ represent the initial hyperparameter values that define the prior distribution. The superscript $s$ denotes the $s^{\text{th}}$ sample out of a total of $S$ samples. Subsequently, we perform all necessary computational steps to derive samples from the target quantities $z_{i}^{(s)}$. Finally, we apply the corresponding elicitation technique to obtain the model-implied statistics $t_{m}=f_{j}(\{z_{i}^{(s)}\})$ from the simulations.
As an example, consider a Binomial regression model with a dummy-coded categorical predictor $x$ with two levels and model parameters $\theta=(\theta_{1},\theta_{2})$ representing the intercept and slope. We decide to query the expert about the quartiles of the expected outcome variable $y$ conditional on each group, $\text{gr}_{1}$ and $\text{gr}_{2}$. That is, we obtain six expert-elicited statistics $Q_{p}(y_{\text{gr}_{i}})$ with $p=0.25,0.50,0.75$ and $i=1,2$. To obtain the corresponding model-implied statistics, we would first sample the model parameters from the initial prior distribution, $\theta^{(s)}\sim p_{\lambda_{0}}(\theta)$ and compute the linear predictor $\eta^{(s)}_{\text{gr}_{i}}=\theta_{1}^{(s)}+\theta_{2}^{(s)}x_{i}$ transformed by an inverse-link function $\sigma$ to obtain the implied event probabilities $p^{(s)}_{\text{gr}_{i}}=\sigma\left(\eta^{(s)}_{\text{gr}_{i}}\right)$. Then we would sample from the likelihood providing us with samples from the implied prior predictive distribution, $y^{(s)}_{\text{gr}_{i}}\sim\text{Binomial}(p^{(s)}_{\text{gr}_{i}},N)$. We repeat this sampling process $S$ times resulting in sets $\{y^{(s)}_{\text{gr}_{i}}\}$ of $S$ samples. Finally, we compute the quartiles from these samples, resulting in the six model-implied statistics $Q_{p}(\{y^{(s)}_{\text{gr}_{i}}\})$ with $p=0.25,0.50,0.75$ and $i=1,2$.
2.4 Discrepancy Evaluation & Training
To evaluate the discrepancy between the model- and expert-elicited statistics, an appropriate loss function is used for each statistic, $L_{m}(t_{m}(\lambda),\hat{t}_{m})$. Different loss functions may be preferable depending on the type of the respective elicited statistic. Since the discrepancy is computed for each statistic within the set of elicited statistics, the total loss comprises multiple loss components, which are combined into a single total loss using a weighted sum:
$L(\lambda)=\sum_{m=1}^{M}\alpha_{m}L_{m}(t_{m}(\lambda),\hat{t}_{m}),$ (1)
where $\alpha_{m}$ denote the weights. The choice of weights $\alpha_{m}$ can be guided by different considerations: One approach is to manually adjust the weights based on the importance of the individual loss components from the expert’s perspective [72]. Alternatively, automated weighting methods, known as loss-balancing methods, can be used to optimally balance the contribution of each individual term to the total gradient [46, 13, 4].
After computing the total loss, $L(\lambda)$, the gradients with respect to the hyperparameters $\lambda$ are calculated. These gradients are then used to adjust the hyperparameters in the opposite direction:
$\lambda^{{}^{\prime}}\leftarrow\lambda-\delta\sum^{M}\alpha_{m}\frac{\partial L_{m}}{\partial\lambda},$ (2)
with $\delta$ being the step size [34]. We employ mini-batch stochastic gradient descent (SGD) with automatic differentiation, facilitated by the (explicit or implicit) reparameterization trick [43, 26]. With the adjusted hyperparameters $\lambda^{\prime}$, new samples from the prior $p_{\lambda^{\prime}}(\theta)$ can be generated and the updated model-implied statistics $t_{m}(\lambda^{\prime})$ computed. We then reassess the discrepancy between the model-implied and expert-elicited statistics and continue to adjust the hyperparameters as needed. This iterative process continues until a convergence criterion is met or a stopping rule is applied.
2.5 Evaluation of Training Results
Convergence Checks
Training is considered successful if the loss in minimized and no further learning occurs. With a proper discrepancy measure as loss function, it is guaranteed that the total loss (and the individual loss components) approach zero as learning progresses. The learning progress can be tracked by examining the loss behavior across epochs, where a decreasing trend is expected until the loss stabilizes at a value close to zero. Further insights into the learning progress can be obtained by tracking the convergence of additional quantities of interest beside the loss, for instance the mean and standard deviation of the marginal priors.
A further method to assist in the examination of convergence is to compute the slope of a linear regression fitted to the loss values over the last $m$ epochs, whereby $m$ should not be taken too large such that the linearity assumption holds. Typically, the overall loss function exhibits exponential decay, but once the loss stabilizes, a linear trend is a good approximation. Ideally, a slope of zero would indicate perfect convergence, though some variation is expected in practice. To evaluate whether the slope remains acceptable, the training algorithm can be run repeatedly with random seeds and the absolute slopes across all replications can be compared. For training runs with the highest final slopes, we can perform a visual inspection of the trajectories of key quantities across epochs, such as the total loss, loss components, and prior means and standard deviations. If these runs show successful convergence, we can reasonably conclude that other runs (with smaller final slopes) have also converged.
Non-uniqueness of Learned Priors
Once the elicited statistics have been learned accurately, we can examine the learned prior distribution that corresponds to this set of expert-elicited statistics. However, there are often multiple optimal values $\lambda^{*}$ that align closely with the elicited statistics yet result in different prior distributions [67, 47, 69]. This lack of uniqueness is anticipated, given the limited information provided by the set of elicited statistics compared to the complexity of the generative model [47]. Furthermore, when target quantities refer solely to the observable space, we only gain access to the model’s overall uncertainty with limited ability to differentiate between aleatoric and epistemic uncertainty [53, 60]. Aleatoric uncertainty, by definition, is irreducible; however, several methods exist to reduce epistemic uncertainty. These include (i) selecting more informative target quantities, (ii) eliciting more detailed expert information (e.g., increasing the number of quantiles in quantile-based elicitation), (iii) adding theory-informed regularization terms to constrain the search space [as illustrated in 47], (iv) improving the initialization method to achieve better starting points, and (v) employing more advanced optimization algorithms when dealing with a large number of model parameters [53].
Model Averaging
Given the (potential) non-uniqueness of the learned prior distributions, it is advisable to run the training multiple times on the same expert data but with different random seeds. This approach helps provide insight into the variability among the learned prior distributions. From the resulting set of candidate priors $p_{\lambda_{r}}(\theta)$ for $r=1,\ldots,R$, we can, in consultation with the domain expert, exclude prior distributions that do not meet the expert’s expectations. For the remaining set of prior distributions, which may be considered practically equivalent from the expert’s perspective, either a single prior distribution can be chosen, or alternatively, model averaging techniques can be applied. In the case of model averaging, the combined prior would take the form:
$p_{\lambda_{1},\ldots,\lambda_{R}}(\theta)=\sum_{r=1}^{R}w_{r}\cdot p_{\lambda_{r}}(\theta)$ (3)
where the weights $w_{r}$ sum to one. The weights $w_{r}$ can either be assigned by the user, based on theoretical assumptions, or selected according to one of several proposed weighting schemes [10]. Here, we apply an approach suggested by Buckland et al [8] for the general purpose of model averaging and selection, where the weights $w_{r}$ are defined as
$w_{r}=\frac{\exp\{-\gamma I_{r}\}}{\sum_{v=1}^{V}\exp\{-\gamma I_{v}\}}.$
In the above expression, $\gamma$ is a scaling factor, and $I$ represents the performance criterion by which the models are weighted. This definition ensures that models with identical performance receive equal weight [8]. In our setting, it is sensible to use the total loss $L(\lambda_{r})$ of each replication $r$ as the performance criterion. Consequently, the “best” model is the one with the smallest total loss, $L(\lambda_{\min})$ [5]. This yields the following definition of weights
$w_{r}=\frac{\exp\{-\gamma\Delta_{r}(L)\}}{\sum_{v=1}^{V}\exp\{-\gamma\Delta_{v}(L)\}}.$ (4)
where $\Delta_{r}(L)=L(\lambda_{r})-L(\lambda_{\min})$. The difference is used primarily for computational reasons to ensure numerical stability [10]. The formulation in Eq. (4) can be interpreted as the probability of replication $r$ representing the best prior, given the collection of learned prior distributions [71]. The scaling factor $\gamma$ controls the influence of each replication’s relative performance $\Delta_{r}(L)$. When $\gamma>1$, discrepancies in model performance are amplified, with better models receiving higher weights and worse models receiving lower weights. As $\gamma\rightarrow\infty$, this weighting would approach a one-hot vector, effectively performing model selection. Conversely, when $\gamma<1$, the performance differences between models are smoothed out, reducing the influence of performance discrepancies.
Simulating From an “Oracle”
The simulation-based framework enables us to evaluate the degree of non-identifiability and determine if further regularization is required, even before querying any data from the expert. An effective approach involves using an oracle, where a pre-defined prior distribution serves as the ground truth. From this “true” prior, we perform forward simulations: sampling first from the prior, then from the likelihood, and carrying out all necessary computational steps to generate a set of model-implied statistics corresponding to the ground truth. This set of simulated statistics can then serve as a proxy for the “expert-elicited” statistics.
Using the above approach allows us to assess the self-consistency of the algorithm, the degree of non-identifiability (i.e., how informative are the elicited statistics), and the suitability of the chosen algorithm parameters used for training NFs with mini-batch SGD. Additionally, we can use this ground truth as a benchmark to assess whether the learned prior distributions approximate it reasonably well, even though non-identifiability may prevent us from exactly recovering the underlying “true” prior distribution.
3 Implementation
In all simulation studies, we use a batch size of $128$. To simulate from an oracle (discussed in Section 4.1), we draw once $S=10,000$ samples from the true joint prior and then compute the corresponding set of “expert”-elicited statistics. During training with mini-batch SGD we draw $S=200$ samples from the joint prior. To evaluate the discrepancy between the model-implied and expert-elicited statistics, we use a biased Maximum Mean Discrepancy (MMD) loss with an energy kernel [36, 25], following our previous success with the metric [6]. For the correlation between parameters, however, we apply a squared-error loss (for details see Section 4.1). The total loss is computed as a weighted sum of the individual loss components, with unit weights assigned to the MMD loss components and a weight of $0.1$ applied to the squared-error loss components.
The architecture of the NFs is rather simple compared to typical applications in deep learning [e.g., 42] and consists of a standard multivariate Gaussian as a base distribution and three affine coupling blocks. Each coupling block consists of two dense layers with $128$ units and ReLU activation functions. In Simulation Study 1, the number of learnable hyperparameters is $\mid\lambda\mid=202,776$, while in the case studies corresponding to Simulation Study 2, it is $\mid\lambda\mid=205,872$. In Simulation Study 1, we assume a discrete likelihood and use the softmax-gumble trick to compute the gradients of a discrete random variable [for more information, see 6]. The gradients of the total loss with respect to the hyperparameters are used to update the hyperparameters $\lambda$. Optimization was performed with the Adam optimizer for $600$ epochs in Simulation Study 1 and $1500$ epochs in Simulation Study 2. We used a fixed learning rate of $0.00025$ for Simulation Study 2, Scenario 1 & 2, and $0.0001$ for Simulation Study 1 and Stimulation Study 2, Scenario 3. These algorithm (hyper-)parameter values were obtained by manual tuning until good performance across all of our case studies was achieved.
All simulations were implemented in Python 3.11.10 using the following libraries: TensorFlow 2.14 [1], TensorFlow Probability 0.22.1 [18], NumPy 1.24.3 [37], pandas 2.2.3 [49], and BayesFlow 1.1.6 [63]. Simulations were run on the Linux HPC cluster at Technical University Dortmund (LiDO3) on high-end CPUs.
4 Simulation Studies
4.1 General Setup
[…]
Model Generative Model Ground Truth
Binomial — M1 $y_{i}\sim\text{Binomial}(p_{i},30)$ $\beta_{0}\sim\text{Normal}(0.1,0.1)$
(Section 4.2) $p_{i}=\text{sigmoid}(\beta_{0}+\beta_{1}x_{i})$ $\beta_{1}\sim\text{Normal}(-0.1,0.3)$
Normal models
Scenario 1 — M2 $y_{i}\sim\text{Normal}(\mu_{i},\sigma)$ $\beta_{0}\sim~{}\text{Normal}(10,2.5)$
 (Section 4.3.1) $\mu_{i}=\beta_{0}+\beta_{1}x_{1,i}+\beta_{2}x_{2,i}$ $\beta_{1}\sim\text{Normal}(7,1.3)$
$\beta_{2}\sim\text{Normal}(2.5,0.8)$
$\sigma\sim\text{Gamma}(5,2)$
[…]
Scenario 2 — M3 see M2 $\beta_{0}\sim~{}\text{Normal}(10,2.5)$
 (Section 4.3.2) $\beta_{1}\sim\text{SkewNormal}(7,1.3,4)$
$\beta_{2}\sim\text{SkewNormal}(2.5,0.8,4)$
$\sigma\sim\text{Gamma}(5,2)$
[…]
Scenario 3 — M4
 (Section 4.3.3) see M2 $\beta\sim\text{Mv-Normal}\left(\begin{bmatrix}10\\ 7\\ 2.5\end{bmatrix},\;\text{D}(s)\,\textbf{R}\,\text{D}(s)\right)$
$\textbf{R}=\begin{bmatrix}1.&0.3&-0.3\\ 0.3&1.&-0.2\\ -0.3&-0.2&1.\end{bmatrix}$
$s=(2.5,1.3,0.8)$
$\sigma\sim\text{Gamma}(5,2)$
Table 2: Overview of generative models and ground truth in the simulation studies. Symbols: $\mathbf{R}$ denotes the correlation matrix and $s$ the vector of standard deviations.
Simulated Expert as Ground Truth
For the simulation studies in this paper, we simulate expert input using synthetic data generated from a predefined joint prior representing the ground truth (i.e., oracle, see Section 2.5 for details). For simplicity, we use a parametric prior with fixed hyperparameter values as ground truth for each case study. Further details on the exact specification of the “true” joint prior is provided in the respective case-study sections and in Table 2.
Selection of Target Quantities and Elicitation Techniques
Following recommendations in the field, we focused on observable quantities when selecting target quantities for our simulation studies (see Section 2.2 for details). For the Binomial regression model, we used the expected observations of the outcome variable conditional on two observations of the continuous predictor, $y\mid x_{0}$ and $y\mid x_{1}$. For the normal regression model, we considered the prior predictive observations conditional on the three groups, $y\mid\text{gr}_{i}$ for $i=1,2,3$.
However, we had to deviate from the recommendation to select observable target quantities in two respects. First, we included $R^{2}$ as an additional target quantity for the normal model (Simulation Study 2), motivated by the following reasoning: if we only have information about the outcome variable for each of the three groups, we provide the learning algorithm with no means to differentiate between parameter uncertainty and data uncertainty. The coefficient of determination, $R^{2}$, represents the ratio of parameter uncertainty to total uncertainty (which encompasses both data and parameter uncertainty). Consequently, it offers more information for the learning algorithm to differentiate between these two types of uncertainty, even though it may be somewhat more challenging for a domain expert to interpret. However, as $R^{2}$ is commonly reported as a metric in a regression context, we would argue that this quantity is still well understood by domain experts.
Second, we included the pairwise correlation between model parameters in the set of target quantities for each case study. If independence between model parameters is assumed, all correlations are set to zero; otherwise, we use the exact correlation structure indicated by the ground truth. We acknowledge that this information cannot be reasonably demanded from a domain expert in most cases. However, we currently lack suitable elicitation methods for asking domain experts about interpretable quantities that provide sufficient information regarding the correlation structure between the model parameters. A future task will be to develop such elicitation methods and to then incorporate these methods into our computational framework.
Regarding the selection of elicitation techniques, we follow the recommendation to query quantiles from a domain expert [40]. For each target quantity, except for the correlation information, we elicit five quantiles: $5\%,25\%,50\%,75\%,95\%$. For the correlation values, we assume that the expert provides a single point estimate corresponding to a moment-based elicitation approach. To assess whether the selected set of elicited statistics is sufficiently informative to determine a prior distribution for the model parameters, we conduct a sensitivity analysis for each case study.
Evaluation of Simulation Results
We run each simulation study 30 times with different random seeds. For each replication, we calculate the slope of the loss trajectory over the last 100 epochs and perform a visual convergence check for the five replications with the highest absolute slopes (i.e., worst cases). In this visual check, we examine the trajectories of the total loss, the loss components related to the elicited statistics, and the mean and standard deviation for each marginal prior across epochs. If this visual inspection indicates satisfactory convergence, we consider the convergence of all other replications to be successful as well. Additionally, we check the match between the final model-implied and expert-elicited statistics. Finally, we examine the learned prior distributions for each replication and compute the model average across the 30 learned prior distributions.
Summary of Analysis Workflow
Table 3 provides a summary of the analysis workflow followed in each of the subsequent simulation studies.
Task Approach Insights
Setup & Elicitation Stage
Select the set of elicited statistics and evaluate its informativeness. (1) Simulate elicited statistics by defining a true joint prior (i.e., the oracle). (2) Conduct a sensitivity analysis. Can the set of elicited statistics provide sufficient information to learn the key aspects of the joint prior?
Fitting Stage
Evaluate variability in learned priors to assess non-identifiability. Run the training algorithm several times with different random seeds. Do different training runs yield different results?
[…]
Evaluation Stage
Examine the learned priors across the different seeds Visually inspect the marginal priors and correlations between model parameters. How much do the learned priors vary between different seeds? Were any degenerate prior distributions learned? Do we observe any form of multi-modality? Can any of the learned priors be ruled out as unreasonable post hoc?
[…]
Table 3: Overview of the analysis workflow used in the simulation studies.
4.2 Simulation Study 1: Binomial Likelihood with Independent Normal Priors
We introduce the general idea of our method using a Binomial model with a logit link (sigmoid response function) and one continuous predictor $x=1,\ldots,50$, scaled by its standard deviation. In the following we refer to this model as M1:
$\begin{split}y_{i}&\sim\text{Binomial}(p_{i},30)\\ p_{i}&=\text{sigmoid}(\beta_{0}+\beta_{1}x_{i})\\ \beta_{0},\beta_{1}&\sim p_{\lambda}(\beta_{0},\beta_{1})\\ \theta&\equiv(\beta_{0},\beta_{1})\end{split}$ (M1)
The goal is to learn a joint prior for the model parameters $\beta_{0}$ (intercept) and $\beta_{1}$ (slope), assuming independence between these parameters.
Setup & Elicitation Stage
As elicited statistics, we select two observations from the continuous predictor $x$, denoted by $x_{0}$ and $x_{1}$, which correspond to the $25\%$ and $75\%$ quantiles of $x_{i}$. We select specific quantiles rather than randomly sampling two observations from the predictor to avoid the observations being too close together, which would reduce informativeness. From the two selected observations, we compute five quantiles from the prior predictive distribution to represent the elicited statistics (see Section 4.1 for details). To obtain the “expert”-elicited statistics, we define a true prior that represents the ground truth and simulate from the generative model in forward mode, computing the corresponding true-elicited statistics. The true joint prior is defined by independent normal distributions for each model parameter: $\beta_{0}\sim\text{Normal}(0.1,0.1)$ and $\beta_{1}\sim\text{Normal}(-0.1,0.3)$.
To assess the informativeness of the selected set of elicited statistics, we conduct a sensitivity analysis, with the results shown in Figure 3. In each row, a single hyperparameter is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line. The columns present the two elicited statistics, specifically the five quantiles (in shades of blue and green) for $y\mid x_{0}$ and $y\mid x_{1}$. The horizontal gray line is included for reference, with the upper two rows indicating results for the intercept parameter $\beta_{0}$ and the lower two rows for the slope parameter $\beta_{1}$. The results of the sensitivity analysis indicate that the selected elicited statistics provide sufficient informativeness for the model hyperparameters.
Figure 3: Binomial model—Sensitivity Analysis: Rows represent hyperparameters of each model parameter. Columns represent elicited statistics: five quantiles for $y\mid x_{0}$ and $y\mid x_{1}$. Quantiles are depicted in different colors. In each row, the corresponding hyperparameter, is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line. The horizontal gray line is included for reference, with the upper two rows indicating results for the intercept parameter $\beta_{0}$ and the lower two rows for the slope parameter $\beta_{1}$.
Fitting Stage
We ran the training algorithm 30 times with different random seeds and conducted an initial convergence check by inspecting the slope of the last 100 epochs for each replication. In Figure 4, the upper row shows the absolute slope (scaled by a factor of 100) for each replication. In the lower row, the learning trajectory of the total loss is shown for the best-performing (leftmost) and the four worst-performing replications. We observe a near-zero slope for the best model, indicating convergence, while the negative slopes for the worst-performing seeds suggest that the learning algorithm continues to make progress. However, a visual inspection of the learning trajectories for the four worst-performing replications, exemplified in Figure 5 for seed 11, indicates satisfactory convergence for all quantities of interest. This includes the total loss, the individual loss components (upper row), as well as the means and standard deviations of the marginal priors (lower row). Therefore, although there appears to be minor ongoing learning progress, the results will only change very minimally. Overall, we conclude that the training process was successful.
Figure 4: Binomial model—Preliminary convergence check for all 30 replications: The upper plot displays the absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis. The replications (i.e., seeds) are arranged in ascending order based on the magnitude of the absolute slope. The best-performing model, with an almost zero slope, is seed 28, while the worst-performing models, indicated by negative slopes, are seeds 3, 14, 15, and 11. The lower plot illustrates the loss trajectories of the best- and worst-performing seeds over the last 200 epochs. The loss updates for the final 100 epochs are highlighted in green for the best model and in blue for the worst models, with a red line segment indicating the respective slope.
Figure 5: Binomial model—Visual convergence check for seed=11: The plots display the updating trajectory across epochs for various quantities of interest: the total loss (upper left), individual loss components (upper right), and the mean and standard deviation of the marginal priors (lower left and lower right, respectively).
The corresponding final learned elicited statistics for each replication are shown in the first two plots in Figure 6. Each true quantile (x-axis) is plotted against its corresponding learned quantile (y-axis) for each replication. Points that align perfectly with the diagonal dashed line represent a perfect match between true and learned quantiles. Points above the diagonal indicate that the learned quantiles are higher than the true ones (and vice versa for lower points). Additionally, in the rightmost plot in Figure 6, we present the true (red) and learned (orange) correlations between the marginal priors. Consistent with successful convergence, the elicited statistics, including the correlations, are accurately learned across all replications.
Figure 6: Binomial model—Learned elicited statistics for all 30 replications: The two leftmost plots show the learned quantiles (y-axis) vs. the true quantiles (x-axis) of each replication for the two target quantities, $y\mid x_{0}$ and $y\mid x_{1}$. The diagonal line serves as a reference, where points lying on the line indicate a perfect match between learned and true quantiles. The rightmost plot displays the learned (orange) vs. true (red) correlations. Given that we assume independence between model parameters in this scenario, the true correlation is zero. Consistent with convergence, all learned elicited statistics closely match the ground truth.
Evaluation Stage
Finally, we can examine the prior distribution learned by each of the 30 replications. Figure 7b displays the learned marginal priors for each replication in grey. Since all priors closely align with the ground truth (indicated by the dotted black line), the different seeds are visually indistinguishable. These results suggest that the set of elicited statistics provides sufficient information to practically identify model M1. This conclusion is further supported by the weights used for model averaging, depicted in Figure 7a. The weights represent the relative performance of each replication based on their total loss (see Section 2.5 for details). All replications perform equally well, resulting in a uniform weighting scheme. The corresponding model average is depicted via a dashed black line in Figure 7b. In this example, all learned marginal priors, the ground truth, and the model average coincide, showcasing an ideal scenario.
Figure 7: Binomial model—Learned marginal priors for all 30 replications: (a) Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss. In this scenario, all replications perform equally well, resulting in uniform weights. (b) The learned marginal priors for each replication are shown in grey. The dotted black line represents the ground truth and the dashed, black line indicates the model average. In this case, the elicited statistics are sufficiently informative to practically identify the generative model, leading the learned priors, the ground truth, and the model average to coincide.
4.3 Simulation Studies 2: Normal Likelihood
In the remaining simulation studies, we aim to evaluate the method’s performance using a model with increased complexity, specifically a normal regression model with standard deviation $\sigma$ and mean $\mu_{i}$, where $\mu_{i}$ depends on a three-level grouping variable encoded via the dummy variables $x_{1}$ and $x_{2}$:
$\begin{split}y_{i}&\sim\text{Normal}(\mu_{i},\sigma)\\ \mu_{i}&=\beta_{0}+\beta_{1}x_{1,i}+\beta_{2}x_{2,i}\\ \beta_{0},\beta_{1},\beta_{2},\sigma&\sim p_{\lambda}(\beta_{0},\beta_{1},\beta_{2},\sigma)\\ \theta&\equiv(\beta_{0},\beta_{1},\beta_{2},\sigma)\end{split}$
with $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$ being the regression coefficients representing the intercept and the two group contrasts, respectively. The goal is to learn a joint prior for the model parameters $\theta$. To systematically assess our method, we use the same model across all upcoming simulation studies, varying only the true joint prior in each scenario:
  • Scenario 1 (M2): Independent normal priors for the regression coefficients are defined as $\beta_{0}\sim~{}\text{Normal}(10,2.5),\beta_{1}\sim\text{Normal}(7,1.3),\beta_{2}\sim\text{Normal}(2.5,0.8)$, with a Gamma prior for the random noise, $\sigma\sim\text{Gamma}(5,2)$
  • Scenario 2 (M3): This setup is identical to M2 but introduces skewed normal priors for $\beta_{1}$ and $\beta_{2}$, where $\beta_{1}\sim\text{SkewNormal}(7,1.3,4)$ [^1] For the skewed normal distribution, we use the implementation from TensorFlow Probability [1], which features the two-piece normal distribution [24]. This distribution is parameterized by location, scale, and shape parameters. The shape parameter controls the skewness, with values above one resulting in a right-skewed distribution. and $\beta_{2}\sim\text{SkewNormal}(2.5,0.8,4)$.
  • Scenario 3 (M4): This scenario introduces a dependency structure among the model coefficients $\beta$ using a correlated multivariate normal distribution, with marginal distributions identical to those in M2. All other aspects remain the same as in M2. The correlated prior is specified as:
[…]
    R $=\begin{bmatrix}1.&0.3&-0.3\\ 0.3&1.&-0.2\\ -0.3&-0.2&1.\end{bmatrix},\quad s=(2.5,1.3,0.8)$
    where $\mathbf{R}$ denotes the correlation matrix and $s$ the vector of standard deviations.
4.3.1 Scenario 1: Independent Normal Priors
Setup & Elicitation Stage
We begin with a brief assessment of the selected elicited statistics, with sensitivity analysis results detailed in Appendix A.2. While the intercept coefficient, $\beta_{0}$, is well-informed by all three target quantities related to the predictive distributions of the groups, the slope coefficients, $\beta_{1}$ and $\beta_{2}$, are each informed by only one target quantity, specifically those corresponding to group 2 and group 3, respectively. Consequently, we anticipate greater difficulty in accurately learning these coefficients. Additionally, we included $R^{2}$ among the target quantities to aid the learning algorithm in distinguishing parameter uncertainty from data uncertainty. However, given the limited data available, we expect considerable variability in the learned model variances.
Fitting Stage
To verify successful convergence, we first examined the absolute slopes across all 30 replications, followed by a detailed visual inspection of the learning trajectories for the loss components and additional quantities of interest in the replications with the highest absolute slopes (see Appendix A.2 for results). Based on this analysis, we conclude that the learning process was successful. The corresponding learned model-implied statistics are shown in Figure 8 and demonstrate a strong match between the learned and true elicited statistics.
Figure 8: Normal model / Scenario 1 (M2) — Learned elicited statistics for all 30 replications: Upper plots show the learned quantiles (y-axis) vs. the true quantiles (x-axis) of each replication for the four target quantities $y\mid\text{gr}_{i}$ with $i=1,2,3$ and $R^{2}$. The diagonal line serves as a reference, where points lying on the line indicate a perfect match between learned and true quantiles. The lower plot displays the learned (orange) vs. true (red) correlations between the model parameters. Given the independence assumption, the true correlation is zero. All learned elicited statistics closely match the ground truth.
Evaluation Stage
The corresponding learned marginal priors of each replication are depicted in Figure 9. Figure 9a displays the distribution of the computed weights used for model averaging. The replication with the highest weight is highlighted in green, while the replication with the lowest weight is highlighted in orange. In Figure 9b, the learned marginal priors of each replication are depicted in grey. The dotted black line refers to the ground truth and the dashed black line to the model average. Additionally, the marginal priors for the best- and worst-performing replications are highlighted in their respective colors. The results suggest that the elicited statistics provide sufficient information to accurately learn the marginal prior of $\beta_{0}$. In contrast, the remaining learned marginal priors display substantial variation. Notably, for $\sigma$, most replications yield a marginal prior concentrated around the mean of a Gamma(5, 2) distribution. This indicates that, while the elicited statistics capture some information about the average level of data variation, they convey limited information about its uncertainty. However, despite the variability in the learned marginal priors, there is no degenerate or otherwise highly outlying distribution whose plausibility we would exclude a priori.
Figure 9: Normal model / Scenario 1 (M2) — Learned marginal priors for all 30 replications: (a) Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss. The highest and lowest weights are highlighted in green and orange, respectively. (b) The learned marginal priors for each replication are shown in grey, with the dotted black line representing the ground truth and the dashed black line indicating the model average. The marginals for the seeds with the highest and lowest weights are highlighted in the respective colors. Although the elicited statistics provide sufficient information to accurately learn the prior for $\beta_{0}$, the substantial variation in the learned priors for the other model parameters indicates a certain lack of model identification.
4.3.2 Scenario 2: Skewed Normal Priors
Setup & Elicitation Stage
This scenario differs from the previous one only in the ground truth, where we introduce a skew-normal marginal prior for the two slope coefficients, $\beta_{1}$ and $\beta_{2}$. However, since we use the same set of ”elicited” statistics, the sensitivity analysis (see Appendix A.3) already suggests that there is likely insufficient information to disentangle the variance and skewness components for these two slope coefficients; therefore, these components are most likely not identified.
Fitting Stage
[…]
Figure 10: Normal model / Scenario 2 (M3) — Learned elicited statistics for all 30 replications: Upper plots show the learned quantiles (y-axis) vs. the true quantiles (x-axis) of each replication for the four target quantities $y\mid\text{gr}_{i}$ with $i=1,2,3$ and $R^{2}$. The diagonal line serves as a reference, where points lying on the line indicate a perfect match between learned and true quantiles. The lower plot displays the learned (orange) vs. true (red) correlations between the model parameters. Given the independence assumption, the true correlation is zero. All learned elicited statistics closely match the ground truth, except for $R^{2}$, where the learned quantiles tend to be overestimated, implying an underestimation of $\sigma$ (see text for details).
Evaluation Stage
Indeed, when examining the learned marginal priors in Figure 10b, we see that for $\sigma$, there is one replication that places a significant amount of prior mass on values near zero. This replication corresponds to the seed with the smallest weight as depicted in Figure 10a and is therefore highlighted in orange. The next step in this case would be to consult with the domain expert on the plausibility of the learned marginal priors, particularly considering whether the marginal prior for $\sigma$, with values close to zero, is meaningful in the context of the study. Since $\sigma$ shares the same scale as the outcome variable $y$, it allows for an assessment of how much or how little noise in the data is considered realistic. This situation illustrates how, even if a domain expert may not have been able to express precise expectations for $\sigma$ initially, the resulting distributions can enable them to retrospectively rule out certain cases they find unrealistic.
Figure 11: Normal model / Scenario 2 (M3) — Learned marginal priors for all 30 replications: (a) Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss. The highest and lowest weights are highlighted in green and orange, respectively. (b) The learned marginal priors for each replication are shown in grey, with the dotted black line representing the ground truth and the dashed black line the model average. The marginals for the seeds with the highest and lowest weights are highlighted in the respective colors. Substantial variation in the learned priors for all model parameters suggests a lack of model identification. However, no degenerate distribution is learned. For $\sigma$, the learned priors exhibit some degree of multi-modality, with one replication in particular placing a significant amount of probability mass on small values of $\sigma$.
4.3.3 Scenario 3: Correlated Normal Priors
Setup & Elicitation Stage
In this final scenario, we introduced a specific dependency structure among the regression coefficients, while all other aspects remained identical to Scenario 1 (M2). The results of the sensitivity analysis, provided in Appendix A.4, indicate that the current set of “elicited” statistics would not provide sufficient information to learn the dependency structure. To address this, we included the pairwise correlations between model parameters as input to the learning algorithm. As a result, the correlation should be accurately learned, given the full information provided.
Fitting Stage
The convergence check indicated successful convergence, with results provided in Appendix A.4. The corresponding match between the model-implied and true “elicited” statistics is shown in Figure 12. We observe a close match for the quantiles of the predictive distribution of each group, while, as noted already in Scenario 2, the quantiles for $R^{2}$ remain more challenging to learn accurately. However, in this scenario, the overestimation of $R^{2}$ is less pronounced, occurring in only one replication, while all other replications appear to learn this statistic accurately. The correlation plot (shown on the bottom of Figure 12) also demonstrates good learning of the dependency information. Therefore, with the provided correlation information, the learning algorithm is able to learn a correlated joint prior.
Figure 12: Normal model / Scenario 3 (M4) — Learned elicited statistics for all 30 replications: Upper plots show the learned quantiles (y-axis) vs. the true quantiles (x-axis) of each replication for the four target quantities $y\mid\text{gr}_{i}$ with $i=1,2,3$ and $R^{2}$. The diagonal line serves as a reference, where points lying on the line indicate a perfect match between learned and true quantiles. The lower plot displays the learned (orange) vs. true (red) correlations between the model parameters. All learned elicited statistics match the ground truth, with some variation observed between replications.
Evaluation Stage
In Figure 13b, the learned marginal priors for each replication are shown, with the computed weights used for model averaging displayed in Figure 13a. As expected, the learned marginal priors for the intercept parameter, $\beta_{0}$, show little variation, while there is substantial variation in the learned marginal priors for the other model parameters, whereby the variation for $\beta_{1}$ is less pronounced compared to $\beta_{2}$. As previously noted through the elicited statistics, there is again one replication that learns a marginal prior for $\sigma$ that places significantly more probability mass on values near zero compared to the learned marginal priors of the other replications. These results again suggest some multi-modality in the learned priors, which could be useful for constraining the search space if certain sets of priors are deemed unrealistic by the domain expert.
Figure 13: Normal model / Scenario 3 (M4) — Learned marginal priors for all 30 replications: (a) Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss. The highest and lowest weights are highlighted in green and orange, respectively. (b) The learned marginal priors for each replication are shown in grey, with the dotted black line representing the ground truth and the dashed black line the model average. The marginals for the seeds with the highest and lowest weights are highlighted in the respective colors.
5 Discussion & Outlook
In this paper, we introduce an elicitation method for learning non-parametric joint priors building on and extending our simulation-based framework [6]. This work demonstrates that our framework is capable of supporting the development of various elicitation methods, in addition to the previously introduced method for learning independent parametric priors. To learn non-parametric joint priors, we utilize normalizing flows [44] to learn flexible joint priors for the model parameters. Through four simulation studies, we demonstrated the successful learning of such priors informed by expert knowledge. Although we employ normalizing flows in this work, other generative models, such as flow matching [45] or diffusion models [9] are also compatible with our framework and can be readily used as a drop-in replacement.
We view the ability to use the same framework for implementing various types of elicitation methods advantageous. By constructing these methods from shared modules (building blocks), our approach not only facilitates direct comparison between elicitation methods but also supports the establishment of benchmarks within the field. Furthermore, a unified framework facilitates the development of a standardized set of diagnostics and analysis tools, addressing a gap that currently exists in this area [50]. These tools can be employed during the elicitation process to specify target quantities, choose appropriate elicitation techniques, and evaluate the learned prior distributions.
Within our framework, expert knowledge can be specified in various ways—ranging from querying model parameters to observable quantities or any derived statistic—tailored to the specific problem and the expertise of the domain expert being queried. This flexibility is achieved through the inclusion of a module that defines both the quantities to be elicited from the expert (cf. target quantities) and the techniques used to query these quantities (cf. elicitation techniques). This design choice provides users with maximum flexibility in specifying the set of quantities to be elicited from the domain expert. However, it also places the responsibility on the user to define a valid and appropriate set of quantities to query from the expert [68]. We would argue that this is in general a desirable property of a method, as it avoids imposing specific modeling choices on the user. Instead of requiring that the problem is framed in a predefined way, the method allows users to adapt it to their specific needs [64, 68, 66]. However, in practice, this flexibility is only effective if good recommendations, guidelines, and diagnostics are available to help users define an appropriate set of quantities to query from the expert [29].
A significant challenge in this context is assessing whether the chosen set of statistics is sufficiently informative for the corresponding model. The goal is not necessarily to identify a unique prior that fully defines the generative model. Instead, the aim is to determine a set of equivalently plausible priors, which is theoretically justified since a domain expert is unlikely to have a single “true” prior in mind [73]. This introduces at least two additional requirements for an elicitation method. First, it should include an efficient implementation that allows the training algorithm to be executed multiple times with different seeds, enabling the assessment of variability in the learned prior distributions. Second, the elicitation method should ideally offer additional diagnostics or advanced techniques, such as model selection or model averaging, to assist the user [23]. In this work, we propose a model averaging method in which the weights calculated for averaging can also serve as diagnostics, helping to identify potential outliers within the set of learned priors. Identifying these edge cases could provide a valuable starting point for evaluating the plausibility of the learned priors in collaboration with a domain expert.
If the set of learned prior distributions includes degenerate cases, this signals the need for additional constraints. These constraints could involve eliciting further or alternative quantities (i.e., elicited statistics) from the expert and/or adding a regularization term to the total loss function to better guide the learning process. To evaluate the informativeness of the elicited statistics, we use a sensitivity analysis as a preliminary evaluation tool [55]. Additionally, we advocate for employing an oracle (i.e., simulating data from a known ground truth) and conducting multiple training replications with varying random seeds. These strategies provide a solid initial check to determine if the learned priors require further constraints. Overall, this discussion highlights the importance of developing comprehensive tutorials and resources to guide users effectively through the elicitation workflow as a key task for future work.
Another important direction for future work is the development of a technique that can learn the dependency structure of a joint prior while simultaneously querying the expert about interpretable quantities. In this paper, we demonstrated that our method can learn a joint prior with the corresponding dependency structure when informed with the pairwise model parameter correlation. However, for practical applicability, the goal is to build on these results and develop a method that infers the correlation indirectly from the expert by asking about interpretable quantities. Recently, Mikkola et al [51] proposed a method that uses pairwise comparisons or ranking techniques to learn a multivariate prior density, which could serve as an interesting avenue for future research.
Other important design choices within our framework that require appropriate default settings include the specification of the discrepancy measure and the selection of the optimization approach, among other aspects. So far, in all simulation studies, we have used the maximum mean discrepancy [MMD, 36] to evaluate the discrepancy between the model-implied and expert-elicited statistics. The only exception is the correlation information, for which we employed an L2 loss. One key advantage of the MMD is its flexibility in handling a wide range of statistics, from simple point estimates to complex histogram data. However, a notable limitation of the MMD is its high (i.e., quadratic) computational cost. In contrast, the L2 loss is computationally efficient but lacks the capability to process histogram data. Despite the higher computational cost, the MMD consistently showed very good performance and robustness across all simulation studies so far. Therefore, it appears to be a solid choice as a default discrepancy measure. Future work should study the performance of different loss functions depending on different elicited statistics, with the goal to develop user guidelines for selecting an appropriate loss function.
Furthermore, in all simulation studies conducted so far, we used mini-batch SGD to optimize the prior network parameters. Given the generally good performance observed, we believe that this optimization method is a solid default choice. However, we aim to explore additional optimization approaches within our framework, such as Bayesian optimization (as employed by Manderson and Goudie [47]). Comparing these approaches in terms of performance and the level of hyperparameter tuning required could offer valuable insights into the conditions under which optimization method might be more advantageous than the other.
Declarations
Acknowledgements
We thank Dmytro Perepolkin and Luna Fazio for their valuable comments and suggestions that greatly improved the manuscript. All authors gratefully acknowledge the computing time provided on the Linux HPC cluster at Technical University Dortmund (LiDO3), partially funded in the course of the Large-Scale Equipment Initiative by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as project 271512359.
Funding
Not applicable.
Competing interests
The authors have no competing interests to declare that are relevant to the content of this article.
Consent to participate
Not applicable.
Consent for publication
Not applicable.
Data availability
The simulation results are available at https://osf.io/xrzh6/.
Materials availability
Not applicable.
Code availability
[…]
Author contribution
Conceptualization: PB, SR, FB, Methodology: PB, FB, Software: FB, Writing (Original Draft): FB, Writing (Review & Editing): SR, PB, Visualization: FB, Supervision: PB, Funding acquisition: PB, SR
Appendix A Simulation Studies
A.1 Training time
The average training time across the 30 replications was
  • Binomial model: 24 minutes and 10 seconds ($\pm$ 33 seconds).
  • Independent normal scenario: 66 minutes and 50 seconds ($\pm$ 5 min: 53 secs).
  • Skewed normal scenario: 74 minutes and 34 seconds ($\pm$ 9 min: 16 secs).
  • Correlated normal scenario: 73 minutes and 18 seconds ($\pm$ 1 min: 59 secs).
A.2 Independent normal scenario
Figure 14: Normal model / Scenario 1 (M2) — Sensitivity Analysis: Rows represent hyperparameters of each model parameter. Columns represent elicited statistics: five quantiles for $y\mid\text{gr}_{i}$ with $i=1,2,3$ and for $R^{2}$. Quantiles are depicted in different colors. In each row, the corresponding hyperparameter, is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line. The horizontal gray line is included as a reference to distinguish between the different model parameters.
Figure 15: Normal model / Scenario 1 (M2) — Preliminary Convergence Check for all 30 replications: The upper plot displays the absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis. The replications (i.e., seeds) are arranged in ascending order based on the magnitude of the absolute slope. The best-performing model, with an almost zero slope, is seed 13, while the worst-performing models, indicated by negative slopes, are seeds 11, 2, 12, and 28. The lower plot illustrates the loss trajectories of the best- and worst-performing seeds over the last 200 epochs. The loss updates for the final 100 epochs are highlighted in green for the best model and in blue for the worst models, with a red line segment indicating the respective slope.
Figure 16: Normal model / Scenario 1 (M2) — Visual convergence check for seed=28: (a) Updating trajectory across epochs for various quantities of interest: the total loss (upper left), individual loss components (upper right), and the mean and standard deviation of the marginal priors (lower left and lower right, respectively). (b) Learned joint prior with marginal distributions along the diagonal and correlation information in the off-diagonal.
A.3 Skewed normal scenario
Figure 17: Normal model / Scenario 2 (M3) — Sensitivity Analysis: Rows represent hyperparameters of each model parameter. Columns represent elicited statistics: five quantiles for $y\mid\text{gr}_{i}$ with $i=1,2,3$ and for $R^{2}$. Quantiles are depicted in different colors. In each row, the corresponding hyperparameter, is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line. The horizontal gray line is included as a reference to distinguish between the different model parameters.
Figure 18: Normal model / Scenario 2 (M3) — Preliminary Convergence Check for all 30 replications: The upper plot displays the absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis. The replications (i.e., seeds) are arranged in ascending order based on the magnitude of the absolute slope. The best-performing model, with an almost zero slope, is seed 21, while the worst-performing models, indicated by negative slopes, are seeds 15, 7, 8, and 27. The lower plot illustrates the loss trajectories of the best- and worst-performing seeds over the last 200 epochs. The loss updates for the final 100 epochs are highlighted in green for the best model and in blue for the worst models, with a red line segment indicating the respective slope.
Figure 19: Normal model / Scenario 2 (M3) — Visual convergence check for seed=27: (a) Updating trajectory across epochs for various quantities of interest: the total loss (upper left), individual loss components (upper right), and the mean and standard deviation of the marginal priors (lower left and lower right, respectively). (b) Learned joint prior with marginal distributions along the diagonal and correlation information in the off-diagonal.
A.4 Correlated normal scenario
Figure 20: Normal model / Scenario 3 (M4) — Sensitivity Analysis: Rows represent hyperparameters of each model parameter. Columns represent elicited statistics: five quantiles for $y\mid\text{gr}_{i}$ with $i=1,2,3$ and for $R^{2}$. Quantiles are depicted in different colors. In each row, the corresponding hyperparameter, is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line. The horizontal gray line is included as a reference to distinguish between the different model parameters.
Figure 21: Normal model / Scenario 3 (M4) — Preliminary Convergence Check for all 30 replications: The upper plot displays the absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis. The replications (i.e., seeds) are arranged in ascending order based on the magnitude of the absolute slope. The best-performing model, with an almost zero slope, is seed 19, while the worst-performing models, indicated by negative slopes, are seeds 14, 5, 27, and 6. The lower plot illustrates the loss trajectories of the best- and worst-performing seeds over the last 200 epochs. The loss updates for the final 100 epochs are highlighted in green for the best model and in blue for the worst models, with a red line segment indicating the respective slope.
Figure 22: Normal model / Scenario 3 (M4) — Visual convergence check for seed=26: (a) Updating trajectory across epochs for various quantities of interest: the total loss (upper left), individual loss components (upper right), and the mean and standard deviation of the marginal priors (lower left and lower right, respectively). (b) Learned joint prior with marginal distributions along the diagonal and correlation information in the off-diagonal.

### v2

@@
 This knowledge can be represented in various forms, but to integrate it into a Bayesian model, we need to translate it into a formal mathematical language that can be expressed as a prior distribution over the model parameters [61, 59, 50, 30].
-A whole field of research, commonly referred to as (expert) prior elicitation, has emerged around the question of how to gather expert knowledge and translate it into appropriate prior distributions [69, 50, 23].
+A distinct field of research, commonly known as (expert) prior elicitation, has developed around the question of how to gather expert knowledge and translate it into appropriate prior distributions.
+This area of study has a long history, dating back to the 1960s [75, 43, 42], and continues to be an active area of research today [71, 52, 24].
 Garthwaite et al [30] identified four key stages in a prior elicitation process:
   1. 1.
-    Setup stage: In this stage, the problem is defined, an expert is selected, and the quantities to be elicited from the expert (referred to as target quantities) are determined;
+    Setup stage: In this stage, the problem is defined, an expert is selected, and the quantities to be elicited from the expert (in this paper referred to as target quantities) are determined;
   2. 2.
@@
 In this context, elicitation methods aim to provide a systematic and formal procedure for deriving prior distributions based on expert-elicited statistics.
-Early elicitation methods primarily tackled the problem by seeking analytical solutions, such as conjugate models or problem-specific transformation functions resulting in highly model-specific prior elicitation methods [see 50, for a recent review].
+Early elicitation methods primarily tackled the problem by seeking analytical solutions, such as conjugate models or problem-specific transformation functions resulting in highly model-specific prior elicitation methods [see 52, for a recent review].
 This model dependence is partly due to the use of direct elicitation techniques, which involve asking experts directly about model parameters [71, 24].
@@
 In both approaches, we employ mini-batch stochastic gradient descent as the optimization method.
-However, our framework supports different optimization methods thanks to its modular design.
 Main Contributions
@@
 2 Methodology
-In this paper, we propose an elicitation method that builds on and extends our recently introduced simulation-based framework [6].
-The overall idea of this framework closely resembles the approach of prior predictive checks [32]: We simulate from the joint model $p(\theta,y)$ and assess how well the resulting prior predictions align with the expert’s expectations.
-If there is a discrepancy between the expert’s expectations and the simulations, the prior specification needs to be adjusted accordingly.
+In this paper, we propose an elicitation method that extends our recently introduced simulation-based framework [6] to support the learning of non-parametric joint priors.
+The overall workflow remains unchanged; the modifications primarily concern how the prior distributions are specified and learned.
+We will first provide a brief overview of the workflow to establish the background, followed by a more detailed discussion of the modification introduced to the original framework, which is the focus of the current paper.
+The general workflow of the framework closely resembles the approach of prior predictive checks [34]: We simulate from the joint model $p(\theta,y)$ and assess how well the resulting prior predictions align with the expert’s expectations.
+If there is a discrepancy between the expert’s expectations and the model simulations, the prior specification needs to be adjusted accordingly.
 Figure 1 provides a graphical representation of the framework and Table 1 offers a summary of the symbols and notation used in the following sections.
@@
 Thus, in the non-parametric approach, $\lambda$ represents the weights of the deep neural networks within the NFs, whereas in the parametric approach, $\lambda$ denotes the parameters of the associated prior distribution families (i.e., the prior hyperparameters).
+A more detailed introduction to NFs will be provided in the upcoming section.
+In summary, the main difference between the parametric approach (proposed by [6]) and the non-parametric approach (focus of the current paper) is that, in the former, the user must specify a parametric prior distribution family for each model parameter, and the corresponding prior hyperparameters are learned via stochastic optimization.
+In the new extension, the user no longer needs to specify a parametric prior for the model parameters.
+Instead, a complex non-parametric prior distribution is assumed.
+This non-parametric joint prior is represented and learned using deep neural networks (specifically, NFs), with the weights learned via stochastic optimization.
+Table 2 summarizes the main steps in the elicitation process and highlights where the extension modifies the existing workflow of our simulation-based framework.
+Furthermore, we provide in this paper additional diagnostics to support users in interpreting the simulation results.
+Elicitation process parametric-prior approach deep-prior approach
+Setup
+Model specify generative model %
+Priors specify parametric prior distribution families assume non-parametric joint prior (specified via NFs)
+Data specify target quantities and elicited statistics %
+(less expert input needed due to stronger regularized priors) (more expert input required (e.g., dependency information) due to flexible prior)
+Elicitation specify elicitation techniques used to query the domain expert %
+Fitting apply mini-batch stochastic gradient descent %
+Evaluation provide diagnostics, visualizations, etc. to evaluate simulation results %
+Table 2: Comparison between the deep-prior and parametric-prior approach on a conceptual level.
+Note: The symbol ‘%’ refers to the same content as provided in parametric-prior approach.
+NFs is an abbreviation for Normalizing Flows (see text for details).
 Normalizing Flows
-To learn a non-parametric joint prior for the model parameters, we use normalizing flows (NFs), a family of deep generative models that offers tractable density evaluation [44].
-NFs transform a simple probability distribution, the base distribution $p(u)$, into a complex target distribution — in our case, the joint prior $p_{\lambda}(\theta)$ — through a series of invertible and differentiable mappings $g_{\lambda}$, parameterized by $\lambda$ [44].
+NFs transform a simple probability distribution, known as the base distribution $p(u)$, into a more complex target distribution, which, in our case, is the joint prior $p_{\lambda}(\theta)$.
+This transformation is achieved through a series of invertible and differentiable mappings $g_{\lambda}$, parameterized by $\lambda$ [46].
 The composition of invertible functions $g_{\lambda}$ yields an explicit form of the density $p_{\lambda}(\theta)$ via the change of variables formula
@@
 Typically, $p(u)$ is chosen to be a simple density, such as a standard Gaussian or a uniform distribution [46].
-Figure 2 depicts the use of NFs for learning a non-parameteric prior distribution within our simulation-based framework.
-Figure 2: Learning non-parametric prior distributions using normalizing flows within our simulation-based framework.
-Normalizing flows transform a simple probability distribution, the base distribution $p(u)$, into a complex target distribution — in our case, the joint prior $p_{\lambda}(\theta)$ — through a series of invertible and differentiable mappings $g_{\lambda}$, parameterized by $\lambda$.
 For NFs to be practical, the mapping $g_{\lambda}$ should be easily invertible, computationally efficient, and sufficiently expressive to model any target distribution of interest [46].
@@
 Elicitation of correlations in the context of joint priors has been studied among others by Gokhale and Press [35], Clemen et al [11], and Dickey et al [17] as reviewed by Mikkola et al [52].
-In our elicitation method, we represent an elicitation technique as a function $f_{j}$ of a target quantity $z_{i}$.
+In our elicitation method, we represent an elicitation technique as a function $f_{j}$ for $j=1,\ldots,J$, of a target quantity $z_{i}$.
 The resulting set of elicited target quantities is referred to as elicited statistics,  $\{t_{m}(\lambda)\}$, where $t_{m}(\lambda)=f_{j}(z_{i})$ and $m$ refers to the corresponding $i\times j$ combination.
-In the following, we sometimes keep the dependency on $\lambda$ implicit and simply write $t_{m}$ instead of $t_{m}(\lambda)$.
+In the following, we sometimes omit the explicit dependence on $\lambda$ and simply write $t_{m}$ instead of $t_{m}(\lambda)$.
+Note also that, depending on the elicitation technique employed, $t_{m}$ may accept either a single value or a set of values as input.
+To illustrate this point more clearly, consider the case in which we elicit expert knowledge about the predictive distribution conditional on a category of a categorical predictor, i.e., $z=y\mid\text{cat}$.
+To elicit this information, we query the expert regarding the $25\%,50\%$, and $75\%$ quantiles of this predictive distribution.
+Thus, we have $t=\{Q_{25\%}(y\mid\text{cat}),\,Q_{50\%}(y\mid\text{cat}),\,Q_{75\%}(y\mid\text{cat})\}$.
 Sensitivity Analysis
@@
 In prior elicitation research, an entire subfield has focused on determining how to conduct optimal expert interviews, taking into account factors such as cognitive biases and heuristics.
-While it is beyond the scope of this paper to review this literature, several comprehensive reviews are available [e.g., 50, 69, 23].
-At this point, we will skip the actual interrogation process and assume that the expert information has been successfully gathered, providing us with a set of expert-elicited statistics, denoted by $\hat{t}_{m}$.
+While it is beyond the scope of this paper to review this literature, several comprehensive reviews are available [e.g., 52, 71, 24].
+At this point, we will skip the actual interrogation process and assume that the expert information has been successfully gathered, providing us with a set of expert-elicited statistics, individually denoted by $\hat{t}_{m}$ and together as $\{\hat{t}_{m}\}$.
 Given the set of expert-elicited statistics, $\{\hat{t}_{m}\}$, we can evaluate the discrepancy between the expert expectations and the model-implied statistics, $\{t_{m}\}$.
@@
 Finally, we apply the corresponding elicitation technique to obtain the model-implied statistics $t_{m}=f_{j}(\{z_{i}^{(s)}\})$ from the simulations.
-As an example, consider a Binomial regression model with a dummy-coded categorical predictor $x$ with two levels and model parameters $\theta=(\theta_{1},\theta_{2})$ representing the intercept and slope.
-We decide to query the expert about the quartiles of the expected outcome variable $y$ conditional on each group, $\text{gr}_{1}$ and $\text{gr}_{2}$.
-That is, we obtain six expert-elicited statistics $Q_{p}(y_{\text{gr}_{i}})$ with $p=0.25,0.50,0.75$ and $i=1,2$.
-To obtain the corresponding model-implied statistics, we would first sample the model parameters from the initial prior distribution, $\theta^{(s)}\sim p_{\lambda_{0}}(\theta)$ and compute the linear predictor $\eta^{(s)}_{\text{gr}_{i}}=\theta_{1}^{(s)}+\theta_{2}^{(s)}x_{i}$ transformed by an inverse-link function $\sigma$ to obtain the implied event probabilities $p^{(s)}_{\text{gr}_{i}}=\sigma\left(\eta^{(s)}_{\text{gr}_{i}}\right)$.
-Then we would sample from the likelihood providing us with samples from the implied prior predictive distribution, $y^{(s)}_{\text{gr}_{i}}\sim\text{Binomial}(p^{(s)}_{\text{gr}_{i}},N)$.
-We repeat this sampling process $S$ times resulting in sets $\{y^{(s)}_{\text{gr}_{i}}\}$ of $S$ samples.
-Finally, we compute the quartiles from these samples, resulting in the six model-implied statistics $Q_{p}(\{y^{(s)}_{\text{gr}_{i}}\})$ with $p=0.25,0.50,0.75$ and $i=1,2$.
 2.4 Discrepancy Evaluation & Training
@@
 where $\alpha_{m}$ denote the weights.
-The choice of weights $\alpha_{m}$ can be guided by different considerations: One approach is to manually adjust the weights based on the importance of the individual loss components from the expert’s perspective [72].
+The choice of weights $\alpha_{m}$ can be guided by different considerations: One approach involves manually adjusting the weights, guided either by the expert’s assessment of the importance of individual loss components or by ensuring that the effects of different magnitudes, arising from varying scales, are balanced, such that no single loss component dominates the others [74].
 Alternatively, automated weighting methods, known as loss-balancing methods, can be used to optimally balance the contribution of each individual term to the total gradient [48, 13, 3].
@@
 Non-uniqueness of Learned Priors
-Once the elicited statistics have been learned accurately, we can examine the learned prior distribution that corresponds to this set of expert-elicited statistics.
-However, there are often multiple optimal values $\lambda^{*}$ that align closely with the elicited statistics yet result in different prior distributions [67, 47, 69].
+Once the elicited statistics have been learned accurately, we can examine the corresponding learned prior distribution.
+As there are often multiple optimal values $\lambda^{*}$ that align closely with the same set of elicited statistics, multiple replications will yield different prior distributions [69, 49, 71].
 This lack of uniqueness is anticipated, given the limited information provided by the set of elicited statistics compared to the complexity of the generative model [49].
@@
 Aleatoric uncertainty, by definition, is irreducible; however, several methods exist to reduce epistemic uncertainty.
-These include (i) selecting more informative target quantities, (ii) eliciting more detailed expert information (e.g., increasing the number of quantiles in quantile-based elicitation), (iii) adding theory-informed regularization terms to constrain the search space [as illustrated in 47], (iv) improving the initialization method to achieve better starting points, and (v) employing more advanced optimization algorithms when dealing with a large number of model parameters [53].
+These include (i) selecting more informative target quantities, (ii) eliciting more detailed expert information (e.g., increasing the number of quantiles in quantile-based elicitation), (iii) adding theory-informed regularization terms to constrain the search space [as illustrated in 49], (iv) improving the initialization method to achieve better starting points, and (v) employing more advanced optimization algorithms when dealing with a large number of model parameters [55].
 Model Averaging
@@
 This set of simulated statistics can then serve as a proxy for the “expert-elicited” statistics.
-Using the above approach allows us to assess the self-consistency of the algorithm, the degree of non-identifiability (i.e., how informative are the elicited statistics), and the suitability of the chosen algorithm parameters used for training NFs with mini-batch SGD.
-Additionally, we can use this ground truth as a benchmark to assess whether the learned prior distributions approximate it reasonably well, even though non-identifiability may prevent us from exactly recovering the underlying “true” prior distribution.
+This oracle approach allows us to assess the self-consistency of the algorithm, the degree of non-identifiability (i.e., how informative are the elicited statistics), and the suitability of the chosen algorithm parameters used for training NFs with mini-batch SGD (such as the learning rate or the number of epochs).
+However, there is one important caveat when using an oracle: the user might be tempted to compare the learned prior distributions against the true prior distributions (which are now available).
+Specifically, if there is an unsatisfactory match between the true and learned prior distributions, the user might interpret this mismatch as a failure in learning.
+It is important to remember, however, that the method only knows about the elicited statistics, which are incorporated into the loss function.
+Thus, when the elicited statistics do not incorporate information about the model parameters but instead pertain to observable quantities, there is, from the method’s perspective, no concept of a true prior.
+In particular, if the true and learned elicited statistics match but there is a mismatch in the learned prior, this is likely a result of an underidentified problem, and the objective function requires additional information to properly constrain the problem.
 3 Implementation
@@
 During training with mini-batch SGD we draw $S=200$ samples from the joint prior.
-To evaluate the discrepancy between the model-implied and expert-elicited statistics, we use a biased Maximum Mean Discrepancy (MMD) loss with an energy kernel [36, 25], following our previous success with the metric [6].
-For the correlation between parameters, however, we apply a squared-error loss (for details see Section 4.1).
-The total loss is computed as a weighted sum of the individual loss components, with unit weights assigned to the MMD loss components and a weight of $0.1$ applied to the squared-error loss components.
-The architecture of the NFs is rather simple compared to typical applications in deep learning [e.g., 42] and consists of a standard multivariate Gaussian as a base distribution and three affine coupling blocks.
+To evaluate the discrepancy between the model-implied and expert-elicited statistics, we use a squared, biased Maximum Mean Discrepancy (MMD) loss with an energy kernel [38, 27], following our previous success with this metric [6].
+For the correlation between parameters, however, we apply a squared-error loss (L2 loss; for details see Section 4.1).
+The total loss is computed as a weighted sum of the individual loss components.
+The values of the respective weights are discussed together with the selection of the elicited statistics whereby each elicited statistic represents one loss component (see Section 4.1).
+The architecture of the NFs is rather simple compared to typical applications in deep learning [e.g., 44] and consists of a standard multivariate Gaussian as a base distribution and three affine coupling blocks.
 Each coupling block consists of two dense layers with $128$ units and ReLU activation functions.
-In Simulation Study 1, the number of learnable hyperparameters is $\mid\lambda\mid=202,776$, while in the case studies corresponding to Simulation Study 2, it is $\mid\lambda\mid=205,872$.
+In Simulation Study 1, the number of learnable hyperparameters is $|\lambda|=202,776$, while in the case studies corresponding to Simulation Study 2, it is $|\lambda|=205,872$.
 In Simulation Study 1, we assume a discrete likelihood and use the softmax-gumble trick to compute the gradients of a discrete random variable [for more information, see 6].
 The gradients of the total loss with respect to the hyperparameters are used to update the hyperparameters $\lambda$.
-Optimization was performed with the Adam optimizer for $600$ epochs in Simulation Study 1 and $1500$ epochs in Simulation Study 2.
+Optimization was performed with the Adam optimizer for $500$ epochs in Simulation Study 1 and $800$ epochs in Simulation Study 2.
 We used a fixed learning rate of $0.00025$ for Simulation Study 2, Scenario 1 & 2, and $0.0001$ for Simulation Study 1 and Stimulation Study 2, Scenario 3.
 These algorithm (hyper-)parameter values were obtained by manual tuning until good performance across all of our case studies was achieved.
-All simulations were implemented in Python 3.11.10 using the following libraries: TensorFlow 2.14 [1], TensorFlow Probability 0.22.1 [18], NumPy 1.24.3 [37], pandas 2.2.3 [49], and BayesFlow 1.1.6 [63].
-Simulations were run on the Linux HPC cluster at Technical University Dortmund (LiDO3) on high-end CPUs.
+All simulations and plots were performed in Python 3.11 using our Python package elicito 0.3.1 [5], which relies, among others, on the following libraries: TensorFlow 2.14 [1], TensorFlow Probability 0.23 [18], NumPy 1.26 [39], pandas 2.2.3 [51], and BayesFlow 1.1.6 [65].
+To ensure the reproducibility of our results, the specific version of elicito used in this paper has been archived on Zenodo and is accessible via the following DOI https://doi.org/10.5281/zenodo.15241853.
+Simulations were executed on a YOGA Pro 9i laptop equipped with GPU acceleration.
 4 Simulation Studies
 4.1 General Setup
-[…]
+In the following, we present four case studies organized across two simulation studies to evaluate the performance of our elicitation method.
+The selection of case studies is motivated by the desire to choose scenarios that are simple and allow us to study the behavior of the method in a well-controlled setting.
+Since this paper presents the extension with NFs for the first time, it is important for us to first develop an understanding of the method’s behavior.
+In future work, we will examine the behavior of our method in more complex models that are closer to real-world applications.
+In Simulation Study 1, we introduce a simple Binomial regression model with one continuous predictor.
+In Simulation Study 2, we present three scenarios based on a normal regression model with a three-level categorical predictor.
+The three normal scenarios differ only in the pre-defined joint prior representing the ground truth.
+Table 3 provides an overview of the generative models and the joint priors representing the ground truth used in the simulation studies.
+Before introducing the simulation studies, we will discuss the decisions made in the simulation setup in greater detail.
+All code and results can be found on GitHub (https://github.com/florence-bockting/non-parametric-prior) and OSF (https://osf.io/xrzh6/), respectively.
+Information about the training time for each simulation study is provided in Appendix A.1.
 Model Generative Model Ground Truth
@@
 $\sigma\sim\text{Gamma}(5,2)$
-Table 2: Overview of generative models and ground truth in the simulation studies. Symbols: $\mathbf{R}$ denotes the correlation matrix and $s$ the vector of standard deviations.
+Table 3: Overview of generative models and ground truth used in each simulation study. Symbols: $\mathbf{R}$ denotes the correlation matrix and $s$ the vector of standard deviations.
 Simulated Expert as Ground Truth
@@
 For simplicity, we use a parametric prior with fixed hyperparameter values as ground truth for each case study.
-Further details on the exact specification of the “true” joint prior is provided in the respective case-study sections and in Table 2.
+Further details on the exact specification of the “true” joint prior is provided in the respective section and in Table 3.
 Selection of Target Quantities and Elicitation Techniques
-Following recommendations in the field, we focused on observable quantities when selecting target quantities for our simulation studies (see Section 2.2 for details).
+Following recommendations in the field, we tried to select mainly observable quantities as target quantities for our simulation studies (see Section 2.2 for details).
 For the Binomial regression model, we used the expected observations of the outcome variable conditional on two observations of the continuous predictor, $y\mid x_{0}$ and $y\mid x_{1}$.
-For the normal regression model, we considered the prior predictive observations conditional on the three groups, $y\mid\text{gr}_{i}$ for $i=1,2,3$.
+For the normal regression model, we considered the prior predictive observations conditional on three groups of the categorical predictor, $y\mid\text{gr}_{i}$ for $i=1,2,3$.
 However, we had to deviate from the recommendation to select observable target quantities in two respects.
@@
 For the correlation values, we assume that the expert provides a single point estimate corresponding to a moment-based elicitation approach.
+In summary, the elicited statistics for Simulation Study 1 consist of five quantiles each for $y\mid x_{0}$ ($t_{1}$) and $y\mid x_{1}$ ($t_{2}$), as well as the correlation information $\rho(\beta_{0},\beta_{1})$ ($t_{3}$).
+Each of these three elicited statistics corresponds to a distinct loss component in the objective function.
+The elicited statistics for Simulation Study 2 consist of five quantiles each for $R^{2}$ ($t_{1}$) and $y\mid\text{gr}_{i}$ with $i=1,2,3$ ($t_{2},t_{3},t_{4}$), as well as the correlation information between the model parameters ($t_{5}$).
+The discrepancy in the correlation values is measured using an L2 loss, whereas the discrepancies for the remaining components are computed using the $\text{MMD}^{2}$ loss.
+To account for differences in scale—arising both from the use of different discrepancy measures and from the varying domains of the elicited statistics—we weight the individual loss components as follows: the correlation loss component is scaled by a factor of $0.1$, the $R^{2}$ loss component by a factor of $10.0$, and all other loss components are uniformly weighted with a factor of $1.0$.
 To assess whether the selected set of elicited statistics is sufficiently informative to determine a prior distribution for the model parameters, we conduct a sensitivity analysis for each case study.
@@
 We run each simulation study 30 times with different random seeds.
-For each replication, we calculate the slope of the loss trajectory over the last 100 epochs and perform a visual convergence check for the five replications with the highest absolute slopes (i.e., worst cases).
-In this visual check, we examine the trajectories of the total loss, the loss components related to the elicited statistics, and the mean and standard deviation for each marginal prior across epochs.
-If this visual inspection indicates satisfactory convergence, we consider the convergence of all other replications to be successful as well.
-Additionally, we check the match between the final model-implied and expert-elicited statistics.
+As preliminary convergence checks, we calculate the slope of the loss trajectory over the last 100 epochs and perform a visual convergence check for the five replications with the highest absolute slopes (i.e., worst cases).
+Additionally, we examine the trajectories of the total loss, the individual loss components related to each elicited statistic, as well as the mean and standard deviation for each marginal prior across epochs.
+Furthermore, to visually assess whether the learning criterion has been accurately captured, we compare the final model-implied statistics with the expert-elicited statistics.
 Finally, we examine the learned prior distributions for each replication and compute the model average across the 30 learned prior distributions.
-Summary of Analysis Workflow
 Table 4 provides a summary of the analysis workflow followed in each of the subsequent simulation studies.
@@
 Table 4: Overview of the analysis workflow used in the simulation studies.
+Additional simulation results
+For each case study, we conduct two additional sets of simulations, which are not part of the main analysis but are included in the supplementary material for completeness and further comparison.
+The supplementary material can be found on GitHub at https://github.com/florence-bockting/non-parametric-prior.
+As these simulations are complementary to the main focus of the paper, we do not report the results in the main text.
+In the first additional set of simulations, we re-ran all case studies using the model parameters as the “elicited statistics” (instead of an indirect quantity derived from the model parameters).
+In this case, the method trains directly on the model parameters, enabling us to assess whether it is, in principle, capable of recovering the true joint prior distribution when complete information is available.
+Thus, this set of simulations serves as a validation set for our method.
+In the second additional set of simulations, we applied the parametric-prior method proposed in Bockting et al [6] to all case studies.
+This allows for a direct comparison between the performance of the proposed non-parametric approach and the previously introduced parametric variant within the same simulation-based framework.
+The only exception is Scenario 3 of Simulation Study 2, where we introduce a multivariate normal prior.
+In this case, a comparison is currently not possible, as the optimization of a full covariance matrix is not yet implemented in elicito [5], but it is planned for a (near) future release.
 4.2 Simulation Study 1: Binomial Likelihood with Independent Normal Priors
-We introduce the general idea of our method using a Binomial model with a logit link (sigmoid response function) and one continuous predictor $x=1,\ldots,50$, scaled by its standard deviation.
+We introduce the general idea of our method using a Binomial model with a logit link (sigmoid response function) and a single continuous predictor, $x$.
+The predictor was generated by first constructing a sequence from 1 to 50, which was then scaled by its standard deviation.
 In the following we refer to this model as M1:
@@
 Setup & Elicitation Stage
-As elicited statistics, we select two observations from the continuous predictor $x$, denoted by $x_{0}$ and $x_{1}$, which correspond to the $25\%$ and $75\%$ quantiles of $x_{i}$.
+As elicited statistics, we select prior predictions for two values of the continuous predictor, $y\mid x_{0}$ and $y\mid x_{1}$.
+The two values of $x$ correspond to the $25\%$ and $75\%$ quantiles of the predictor.
 We select specific quantiles rather than randomly sampling two observations from the predictor to avoid the observations being too close together, which would reduce informativeness.
-From the two selected observations, we compute five quantiles from the prior predictive distribution to represent the elicited statistics (see Section 4.1 for details).
+From the distribution of the two selected observations, we compute five quantiles which represent the elicited statistics (see Section 4.1 for details).
+Note that $y\mid x_{0}$ and $y\mid x_{1}$ each have an associated distribution, resulting from simulating $y^{(s)}$ for $S$ samples drawn from the joint prior and conditioning on the respective $x$ value.
+The distribution over $y\mid x$ encodes the parameter uncertainty.
 To obtain the “expert”-elicited statistics, we define a true prior that represents the ground truth and simulate from the generative model in forward mode, computing the corresponding true-elicited statistics.
@@
 The columns present the two elicited statistics, specifically the five quantiles (in shades of blue and green) for $y\mid x_{0}$ and $y\mid x_{1}$.
-The horizontal gray line is included for reference, with the upper two rows indicating results for the intercept parameter $\beta_{0}$ and the lower two rows for the slope parameter $\beta_{1}$.
+The upper two rows show results for the intercept parameter $\beta_{0}$ and the lower two rows for the slope parameter $\beta_{1}$.
 The results of the sensitivity analysis indicate that the selected elicited statistics provide sufficient informativeness for the model hyperparameters.
@@
 In each row, the corresponding hyperparameter, is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line.
-The horizontal gray line is included for reference, with the upper two rows indicating results for the intercept parameter $\beta_{0}$ and the lower two rows for the slope parameter $\beta_{1}$.
+The upper two rows indicate results for the intercept parameter $\beta_{0}$ and the lower two rows for the slope parameter $\beta_{1}$.
 Fitting Stage
@@
 We observe a near-zero slope for the best model, indicating convergence, while the negative slopes for the worst-performing seeds suggest that the learning algorithm continues to make progress.
-However, a visual inspection of the learning trajectories for the four worst-performing replications, exemplified in Figure 5 for seed 11, indicates satisfactory convergence for all quantities of interest.
-This includes the total loss, the individual loss components (upper row), as well as the means and standard deviations of the marginal priors (lower row).
-Therefore, although there appears to be minor ongoing learning progress, the results will only change very minimally.
+However, a visual inspection of the learning trajectories of the total loss and individual loss components (Figure 5) and the means and standard deviations of the marginal priors (Figure 5), indicates satisfactory convergence.
+Therefore, although the slope analysis indicates minor ongoing learning progress, no substantial changes in the results are expected if the model is trained for additional epochs.
 Overall, we conclude that the training process was successful.
-Figure 4: Binomial model—Preliminary convergence check for all 30 replications: The upper plot displays the absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis.
-The replications (i.e., seeds) are arranged in ascending order based on the magnitude of the absolute slope.
-The best-performing model, with an almost zero slope, is seed 28, while the worst-performing models, indicated by negative slopes, are seeds 3, 14, 15, and 11.
-The lower plot illustrates the loss trajectories of the best- and worst-performing seeds over the last 200 epochs.
-The loss updates for the final 100 epochs are highlighted in green for the best model and in blue for the worst models, with a red line segment indicating the respective slope.
-Figure 5: Binomial model—Visual convergence check for seed=11: The plots display the updating trajectory across epochs for various quantities of interest: the total loss (upper left), individual loss components (upper right), and the mean and standard deviation of the marginal priors (lower left and lower right, respectively).
-The corresponding final learned elicited statistics for each replication are shown in the first two plots in Figure 6.
-Each true quantile (x-axis) is plotted against its corresponding learned quantile (y-axis) for each replication.
+Figure 3: Binomial model—Preliminary convergence check for all 30 replications: Upper plot: Absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis.
+Replications (i.e., seeds) are arranged in ascending order based on the magnitude of the absolute slope.
+Best-performing model is seed 25, while the worst-performing models are seeds 18, 11, 5, and 0.
+Lower plot: Loss trajectories of the best- and worst-performing seeds over the last 200 epochs.
+Figure 4: Binomial model—Visual convergence check 1: Updating trajectory across epochs for all 30 replications.
+On the left, trajectory of the total loss and on the right, trajectory of the weighted individual loss components.
+Figure 5: Binomial model—Visual convergence check 2: Updating trajectory across epochs for for all 30 replications.
+Trajectory of the mean (left plot) and standard deviation (right plot) of the marginal priors.
+The corresponding final learned elicited statistics for each replication are shown in Figure 6.
+In the first two plots, each true quantile (x-axis) is plotted against its corresponding learned quantile (y-axis) for each replication.
 Points that align perfectly with the diagonal dashed line represent a perfect match between true and learned quantiles.
 Points above the diagonal indicate that the learned quantiles are higher than the true ones (and vice versa for lower points).
-Additionally, in the rightmost plot in Figure 6, we present the true (red) and learned (orange) correlations between the marginal priors.
-Consistent with successful convergence, the elicited statistics, including the correlations, are accurately learned across all replications.
+Accuracy of the learned correlation information is presented in Figure 6.
+The true (red) correlation matches perfectly with the learned (black) correlations.
 Figure 6: Binomial model—Learned elicited statistics for all 30 replications: The two leftmost plots show the learned quantiles (y-axis) vs.
@@
 The diagonal line serves as a reference, where points lying on the line indicate a perfect match between learned and true quantiles.
-The rightmost plot displays the learned (orange) vs.
+The rightmost plot displays the learned (black) vs.
 true (red) correlations.
 Given that we assume independence between model parameters in this scenario, the true correlation is zero.
-Consistent with convergence, all learned elicited statistics closely match the ground truth.
 Evaluation Stage
-Finally, we can examine the prior distribution learned by each of the 30 replications.
-Figure 7b displays the learned marginal priors for each replication in grey.
-Since all priors closely align with the ground truth (indicated by the dotted black line), the different seeds are visually indistinguishable.
-These results suggest that the set of elicited statistics provides sufficient information to practically identify model M1.
-This conclusion is further supported by the weights used for model averaging, depicted in Figure 7a.
-The weights represent the relative performance of each replication based on their total loss (see Section 2.5 for details).
-All replications perform equally well, resulting in a uniform weighting scheme.
-The corresponding model average is depicted via a dashed black line in Figure 7b.
-In this example, all learned marginal priors, the ground truth, and the model average coincide, showcasing an ideal scenario.
-Figure 7: Binomial model—Learned marginal priors for all 30 replications: (a) Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss.
-In this scenario, all replications perform equally well, resulting in uniform weights.
-(b) The learned marginal priors for each replication are shown in grey.
-The dotted black line represents the ground truth and the dashed, black line indicates the model average.
-In this case, the elicited statistics are sufficiently informative to practically identify the generative model, leading the learned priors, the ground truth, and the model average to coincide.
+Finally, we examine the learned prior distributions.
+Figure 7 shows the learned marginal priors for each replication, alongside the results from the model averaging approach.
+The weights used for model averaging are shown in the upper plot of Figure 7 and reflect the relative performance of each replication based on their total loss (see Section 4.1 for details).
+In this case, all replications perform similarly well, resulting in an almost uniform weighting scheme.
+The resulting averaged prior is represented by the solid red line in the lower plot, shown together with the learned priors from each replication.
+Figure 7: Binomial model—Learned marginal priors for all 30 replications: Upper plot: Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss.
+In this scenario, all replications perform equally well, resulting in almost uniform weights.
+Lower plot: The learned marginal priors for each replication and in red the prior average.
+In this case, the elicited statistics are sufficiently informative to effectively identify the generative model, resulting in only minor variations in the learned priors across replications.
 4.3 Simulation Studies 2: Normal Likelihood
-In the remaining simulation studies, we aim to evaluate the method’s performance using a model with increased complexity, specifically a normal regression model with standard deviation $\sigma$ and mean $\mu_{i}$, where $\mu_{i}$ depends on a three-level grouping variable encoded via the dummy variables $x_{1}$ and $x_{2}$:
+In the remaining simulation studies, we aim to evaluate the method’s performance using a model with slightly increased complexity, specifically a normal regression model with standard deviation $\sigma$ and mean $\mu_{i}$, where $\mu_{i}$ depends on a three-level grouping variable encoded via the dummy variables $x_{1}$ and $x_{2}$:
 $\begin{split}y_{i}&\sim\text{Normal}(\mu_{i},\sigma)\\ \mu_{i}&=\beta_{0}+\beta_{1}x_{1,i}+\beta_{2}x_{2,i}\\ \beta_{0},\beta_{1},\beta_{2},\sigma&\sim p_{\lambda}(\beta_{0},\beta_{1},\beta_{2},\sigma)\\ \theta&\equiv(\beta_{0},\beta_{1},\beta_{2},\sigma)\end{split}$
@@
   • Scenario 1 (M2): Independent normal priors for the regression coefficients are defined as $\beta_{0}\sim~{}\text{Normal}(10,2.5),\beta_{1}\sim\text{Normal}(7,1.3),\beta_{2}\sim\text{Normal}(2.5,0.8)$, with a Gamma prior for the random noise, $\sigma\sim\text{Gamma}(5,2)$
-  • Scenario 2 (M3): This setup is identical to M2 but introduces skewed normal priors for $\beta_{1}$ and $\beta_{2}$, where $\beta_{1}\sim\text{SkewNormal}(7,1.3,4)$ [^1] For the skewed normal distribution, we use the implementation from TensorFlow Probability [1], which features the two-piece normal distribution [24].
+  • Scenario 2 (M3): Identical to M2 but introduces skewed normal priors for $\beta_{1}$ and $\beta_{2}$, where $\beta_{1}\sim\text{SkewNormal}(7,1.3,4)$ [^1] For the skewed normal distribution, we use the implementation from TensorFlow Probability [1], which features the two-piece normal distribution [26].
 This distribution is parameterized by location, scale, and shape parameters.
@@
 and $\beta_{2}\sim\text{SkewNormal}(2.5,0.8,4)$.
-  • Scenario 3 (M4): This scenario introduces a dependency structure among the model coefficients $\beta$ using a correlated multivariate normal distribution, with marginal distributions identical to those in M2.
-All other aspects remain the same as in M2.
+  • Scenario 3 (M4): Identical to M2 but introduces a dependency structure among the model coefficients $\beta$ using a correlated multivariate normal distribution, with marginal distributions identical to those in M2.
 The correlated prior is specified as:
@@
 Additionally, we included $R^{2}$ among the target quantities to aid the learning algorithm in distinguishing parameter uncertainty from data uncertainty.
-However, given the limited data available, we expect considerable variability in the learned model variances.
 Fitting Stage
-To verify successful convergence, we first examined the absolute slopes across all 30 replications, followed by a detailed visual inspection of the learning trajectories for the loss components and additional quantities of interest in the replications with the highest absolute slopes (see Appendix A.2 for results).
-Based on this analysis, we conclude that the learning process was successful.
-The corresponding learned model-implied statistics are shown in Figure 8 and demonstrate a strong match between the learned and true elicited statistics.
-Figure 8: Normal model / Scenario 1 (M2) — Learned elicited statistics for all 30 replications: Upper plots show the learned quantiles (y-axis) vs.
+To verify successful convergence, we first examined the absolute slopes across all 30 replications, followed by a detailed visual inspection of the learning trajectories for the loss components and additional quantities of interest (see Appendix A.2 for results).
+The analysis indicates that the learning process was successful.
+The model-implied statistics are shown in Figure 8 and demonstrate a strong match between the learned and true elicited statistics, supporting the assumption of successful convergence.
+Figure 8: Normal model / Scenario 1 (M2) — Learned elicited statistics for all 30 replications: First three plots (from the left) show the learned quantiles (y-axis) vs.
 the true quantiles (x-axis) of each replication for the four target quantities $y\mid\text{gr}_{i}$ with $i=1,2,3$ and $R^{2}$.
 The diagonal line serves as a reference, where points lying on the line indicate a perfect match between learned and true quantiles.
-The lower plot displays the learned (orange) vs.
+The rightmost plot displays the learned (black) vs.
 true (red) correlations between the model parameters.
 Given the independence assumption, the true correlation is zero.
-All learned elicited statistics closely match the ground truth.
 Evaluation Stage
-The corresponding learned marginal priors of each replication are depicted in Figure 9.
-Figure 9a displays the distribution of the computed weights used for model averaging.
-The replication with the highest weight is highlighted in green, while the replication with the lowest weight is highlighted in orange.
-In Figure 9b, the learned marginal priors of each replication are depicted in grey.
-The dotted black line refers to the ground truth and the dashed black line to the model average.
-Additionally, the marginal priors for the best- and worst-performing replications are highlighted in their respective colors.
-The results suggest that the elicited statistics provide sufficient information to accurately learn the marginal prior of $\beta_{0}$.
-In contrast, the remaining learned marginal priors display substantial variation.
-Notably, for $\sigma$, most replications yield a marginal prior concentrated around the mean of a Gamma(5, 2) distribution.
-This indicates that, while the elicited statistics capture some information about the average level of data variation, they convey limited information about its uncertainty.
-However, despite the variability in the learned marginal priors, there is no degenerate or otherwise highly outlying distribution whose plausibility we would exclude a priori.
-Figure 9: Normal model / Scenario 1 (M2) — Learned marginal priors for all 30 replications: (a) Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss.
-The highest and lowest weights are highlighted in green and orange, respectively.
-(b) The learned marginal priors for each replication are shown in grey, with the dotted black line representing the ground truth and the dashed black line indicating the model average.
-The marginals for the seeds with the highest and lowest weights are highlighted in the respective colors.
-Although the elicited statistics provide sufficient information to accurately learn the prior for $\beta_{0}$, the substantial variation in the learned priors for the other model parameters indicates a certain lack of model identification.
+The corresponding learned marginal priors of each replication alongside with the model averaging results are depicted in Figure 9.
+In this scenario, we observe noticeably greater variation in the learned priors across different replications, particularly for $\beta_{1}$ and $\beta_{2}$.
+This variation is also reflected in the model averaging weights, which now deviate clearly from the uniform distribution observed in Simulation Study 1.
+These results suggest that while the elicited statistics provide sufficient information to learn the marginal priors of $\beta_{0}$ and $\sigma$, they are insufficient to uniquely identify the priors for $\beta_{1}$ and $\beta_{2}$.
+It is important to emphasize that the only information available to the method for learning the prior distributions is based on the elicited statistics shown in Figure 8.
+Given this limited set of information, learning a non-parametric joint prior becomes a highly underconstrained problem, where multiple prior distributions can be consistent with the same set of elicited statistics.
+However, this should not be interpreted as a limitation of the method itself, but rather as a consequence of the limited information available to the method for learning the joint prior.
+Figure 9: Normal model / Scenario 1 (M2) — Learned marginal priors across replications: Upper plot: Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss.
+Lower plot: The learned marginal priors for each replication (in black) and the prior average (in red).
 4.3.2 Scenario 2: Skewed Normal Priors
 Setup & Elicitation Stage
-This scenario differs from the previous one only in the ground truth, where we introduce a skew-normal marginal prior for the two slope coefficients, $\beta_{1}$ and $\beta_{2}$.
-However, since we use the same set of ”elicited” statistics, the sensitivity analysis (see Appendix A.3) already suggests that there is likely insufficient information to disentangle the variance and skewness components for these two slope coefficients; therefore, these components are most likely not identified.
-Fitting Stage
-[…]
-Figure 10: Normal model / Scenario 2 (M3) — Learned elicited statistics for all 30 replications: Upper plots show the learned quantiles (y-axis) vs.
+Scenario 2 differs from the previous one only in the ground truth, where we introduce a skew-normal marginal prior for the two slope coefficients, $\beta_{1}$ and $\beta_{2}$.
+In this scenario, we examine whether the method can learn this additional skewness property using the same “elicited statistics” as those introduced in the previous case study.
+The sensitivity analysis (see Appendix A.3) already suggests that there is likely insufficient information to disentangle the variance and skewness components for the two slope coefficients; therefore, we except to observe more variation in the learned prior distributions.
+Fitting & Evaluation Stage
+The convergence diagnostics are depicted in Appendix A.3 and indicate successful convergence.
+The corresponding match between the model-implied and true “elicited” statistics is shown in Figure 10, demonstrating overall good performance.
+Figure 10: Normal model / Scenario 2 (M3) — Learned elicited statistics for all 30 replications: First four plots from the left depict the learned quantiles (y-axis) vs.
 the true quantiles (x-axis) of each replication for the four target quantities $y\mid\text{gr}_{i}$ with $i=1,2,3$ and $R^{2}$.
 The diagonal line serves as a reference, where points lying on the line indicate a perfect match between learned and true quantiles.
-The lower plot displays the learned (orange) vs.
+The rightmost plot displays the learned (black) vs.
 true (red) correlations between the model parameters.
 Given the independence assumption, the true correlation is zero.
-All learned elicited statistics closely match the ground truth, except for $R^{2}$, where the learned quantiles tend to be overestimated, implying an underestimation of $\sigma$ (see text for details).
-Evaluation Stage
-Indeed, when examining the learned marginal priors in Figure 10b, we see that for $\sigma$, there is one replication that places a significant amount of prior mass on values near zero.
-This replication corresponds to the seed with the smallest weight as depicted in Figure 10a and is therefore highlighted in orange.
-The next step in this case would be to consult with the domain expert on the plausibility of the learned marginal priors, particularly considering whether the marginal prior for $\sigma$, with values close to zero, is meaningful in the context of the study.
-Since $\sigma$ shares the same scale as the outcome variable $y$, it allows for an assessment of how much or how little noise in the data is considered realistic.
-This situation illustrates how, even if a domain expert may not have been able to express precise expectations for $\sigma$ initially, the resulting distributions can enable them to retrospectively rule out certain cases they find unrealistic.
-Figure 11: Normal model / Scenario 2 (M3) — Learned marginal priors for all 30 replications: (a) Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss.
-The highest and lowest weights are highlighted in green and orange, respectively.
-(b) The learned marginal priors for each replication are shown in grey, with the dotted black line representing the ground truth and the dashed black line the model average.
-The marginals for the seeds with the highest and lowest weights are highlighted in the respective colors.
-Substantial variation in the learned priors for all model parameters suggests a lack of model identification.
-However, no degenerate distribution is learned.
-For $\sigma$, the learned priors exhibit some degree of multi-modality, with one replication in particular placing a significant amount of probability mass on small values of $\sigma$.
+Figure 11 shows the corresponding learned marginal priors together with the results from prior averaging.
+The results indicate relatively stable learning of the prior distributions across all replications.
+As before, the priors for the slope coefficients exhibit higher variation.
+Most importantly, however, the method is able to learn the introduced skewness.
+Figure 11: Normal model / Scenario 2 (M3) — Learned marginal priors for all 30 replications: Upper plot: Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss.
+Lower plot: The learned marginal priors for each replication (in black) and the prior average (in red).
 4.3.3 Scenario 3: Correlated Normal Priors
-Setup & Elicitation Stage
-In this final scenario, we introduced a specific dependency structure among the regression coefficients, while all other aspects remained identical to Scenario 1 (M2).
-The results of the sensitivity analysis, provided in Appendix A.4, indicate that the current set of “elicited” statistics would not provide sufficient information to learn the dependency structure.
-To address this, we included the pairwise correlations between model parameters as input to the learning algorithm.
-As a result, the correlation should be accurately learned, given the full information provided.
-Fitting Stage
-The convergence check indicated successful convergence, with results provided in Appendix A.4.
-The corresponding match between the model-implied and true “elicited” statistics is shown in Figure 12.
-We observe a close match for the quantiles of the predictive distribution of each group, while, as noted already in Scenario 2, the quantiles for $R^{2}$ remain more challenging to learn accurately.
-However, in this scenario, the overestimation of $R^{2}$ is less pronounced, occurring in only one replication, while all other replications appear to learn this statistic accurately.
-The correlation plot (shown on the bottom of Figure 12) also demonstrates good learning of the dependency information.
-Therefore, with the provided correlation information, the learning algorithm is able to learn a correlated joint prior.
-Figure 12: Normal model / Scenario 3 (M4) — Learned elicited statistics for all 30 replications: Upper plots show the learned quantiles (y-axis) vs.
+In this final scenario, we introduced a dependency structure among the regression coefficients, while all other aspects remained identical to Scenario 1 (M1).
+The results of the sensitivity analysis and the convergence checks are provided in Appendix A.4 and show successful convergence.
+The model-implied and true “elicited” statistics are shown in Figure 12 and indicate a close match for all elicited statistics.
+Figure 12: Normal model / Scenario 3 (M4) — Learned elicited statistics for all 30 replications: First four plots show the learned quantiles (y-axis) vs.
 the true quantiles (x-axis) of each replication for the four target quantities $y\mid\text{gr}_{i}$ with $i=1,2,3$ and $R^{2}$.
 The diagonal line serves as a reference, where points lying on the line indicate a perfect match between learned and true quantiles.
-The lower plot displays the learned (orange) vs.
+The rightmost plot displays the learned (black) vs.
 true (red) correlations between the model parameters.
-All learned elicited statistics match the ground truth, with some variation observed between replications.
-Evaluation Stage
-In Figure 13b, the learned marginal priors for each replication are shown, with the computed weights used for model averaging displayed in Figure 13a.
-As expected, the learned marginal priors for the intercept parameter, $\beta_{0}$, show little variation, while there is substantial variation in the learned marginal priors for the other model parameters, whereby the variation for $\beta_{1}$ is less pronounced compared to $\beta_{2}$.
-As previously noted through the elicited statistics, there is again one replication that learns a marginal prior for $\sigma$ that places significantly more probability mass on values near zero compared to the learned marginal priors of the other replications.
-These results again suggest some multi-modality in the learned priors, which could be useful for constraining the search space if certain sets of priors are deemed unrealistic by the domain expert.
-Figure 13: Normal model / Scenario 3 (M4) — Learned marginal priors for all 30 replications: (a) Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss.
-The highest and lowest weights are highlighted in green and orange, respectively.
-(b) The learned marginal priors for each replication are shown in grey, with the dotted black line representing the ground truth and the dashed black line the model average.
-The marginals for the seeds with the highest and lowest weights are highlighted in the respective colors.
+In Figure 13, the final learned marginal priors for each replication are shown alongside the results from prior averaging.
+Results show substantial variation in the learned priors for $\beta_{2}$, whereas the priors for all other parameters show relatively high consistency across replications.
+Since the focus of this case study is specifically on the correlation structure of the regression coefficients, it is informative to also examine the joint prior distribution, which is shown in Figure 14.
+The joint priors from different replications are shown in different colors.
+Although there is clear variation in the learned joint priors, the method is able to capture the underlying correlational pattern.
+Although this result is not particularly surprising, as we provided the method with the exact correlation information between the model parameters, it nonetheless indicates that the method is capable of learning the correlational structure of a joint prior when provided with the relevant information.
+Therefore, future work is needed to develop elicitation methods that enable the method to obtain relevant information by querying domain experts indirectly.
+Figure 13: Normal model / Scenario 3 (M4) — Learned marginal priors for all 30 replications: Upper plot: Computed weights per replication used for model averaging, reflecting the relative performance of each model based on the final total loss.
+Lower plot: The learned marginal priors for each replication (black) and the prior average (red).
+Figure 14: Normal model / Scenario 3 (M4) — Learned joint prior for all 30 replications Each replication is shown in a different color.
+The correlation structure of the regression coefficients used for learning the prior is $\rho(\beta_{0},\beta_{1})=0.3,\rho(\beta_{1},\beta_{2})=-0.2,$ and $\rho(\beta_{0},\beta_{2})=-0.3$.
+Independence between $\sigma$ and correlation coefficients is assumed.
 5 Discussion & Outlook
-In this paper, we introduce an elicitation method for learning non-parametric joint priors building on and extending our simulation-based framework [6].
-This work demonstrates that our framework is capable of supporting the development of various elicitation methods, in addition to the previously introduced method for learning independent parametric priors.
-To learn non-parametric joint priors, we utilize normalizing flows [44] to learn flexible joint priors for the model parameters.
-Through four simulation studies, we demonstrated the successful learning of such priors informed by expert knowledge.
+In this paper, we introduce an extension of our simulation-based framework [6] for learning non-parametric joint priors.
+This extension builds on using normalizing flows [46] to learn flexible joint priors for the model parameters.
 Although we employ normalizing flows in this work, other generative models, such as flow matching [47] or diffusion models [9] are also compatible with our framework and can be readily used as a drop-in replacement.
-We view the ability to use the same framework for implementing various types of elicitation methods advantageous.
-By constructing these methods from shared modules (building blocks), our approach not only facilitates direct comparison between elicitation methods but also supports the establishment of benchmarks within the field.
-Furthermore, a unified framework facilitates the development of a standardized set of diagnostics and analysis tools, addressing a gap that currently exists in this area [50].
-These tools can be employed during the elicitation process to specify target quantities, choose appropriate elicitation techniques, and evaluate the learned prior distributions.
-Within our framework, expert knowledge can be specified in various ways—ranging from querying model parameters to observable quantities or any derived statistic—tailored to the specific problem and the expertise of the domain expert being queried.
-This flexibility is achieved through the inclusion of a module that defines both the quantities to be elicited from the expert (cf.
-target quantities) and the techniques used to query these quantities (cf.
-elicitation techniques).
+Across four simulation studies, we demonstrated successful learning of joint priors based primarily on interpretable quantities that can be meaningfully provided by domain experts.
+One exception is the use of correlation information, which the current method requires to reasonably constrain the optimization problem.
+Future work is required to integrate in our method an elicitation technique that infers the correlation indirectly from the expert by asking about interpretable quantities.
+Recently, Mikkola et al [53] proposed a method that uses pairwise comparisons or ranking techniques to learn a multivariate prior density, which could serve as an interesting avenue for future research.
+With the extension of our originally proposed simulation-based framework to support non-parametric prior distributions, we offer an elicitation framework, implemented in elicito [5], which can accommodate various types of elicitation methods within a single, unified structure.
+This facilitates the development of a standardized set of diagnostics and analysis tools, addressing a current gap in the literature [52].
+In addition to accommodating different types of prior distributions, another strength of the framework is that expert knowledge can be specified in various ways.
+This ranges from querying model parameters to observable quantities or any derived statistics, allowing the queried information to be tailored to the specific problem and the expertise of the domain expert.
 This design choice provides users with maximum flexibility in specifying the set of quantities to be elicited from the domain expert.
 However, it also places the responsibility on the user to define a valid and appropriate set of quantities to query from the expert [70].
-We would argue that this is in general a desirable property of a method, as it avoids imposing specific modeling choices on the user.
-Instead of requiring that the problem is framed in a predefined way, the method allows users to adapt it to their specific needs [64, 68, 66].
+We would argue that this flexibility is in general a desirable property of a method, as it avoids imposing specific modeling choices on the user.
+Thus, instead of requiring that the problem is framed in a predefined way, the method allows users to adapt it to their specific needs [66, 70, 68].
 However, in practice, this flexibility is only effective if good recommendations, guidelines, and diagnostics are available to help users define an appropriate set of quantities to query from the expert [31].
@@
 The goal is not necessarily to identify a unique prior that fully defines the generative model.
-Instead, the aim is to determine a set of equivalently plausible priors, which is theoretically justified since a domain expert is unlikely to have a single “true” prior in mind [73].
+Instead, the aim is to determine a set of equivalently plausible priors, supporting the assumption that a domain expert is unlikely to have a single “true” prior in mind [75].
 This introduces at least two additional requirements for an elicitation method.
@@
 Overall, this discussion highlights the importance of developing comprehensive tutorials and resources to guide users effectively through the elicitation workflow as a key task for future work.
-Another important direction for future work is the development of a technique that can learn the dependency structure of a joint prior while simultaneously querying the expert about interpretable quantities.
-In this paper, we demonstrated that our method can learn a joint prior with the corresponding dependency structure when informed with the pairwise model parameter correlation.
-However, for practical applicability, the goal is to build on these results and develop a method that infers the correlation indirectly from the expert by asking about interpretable quantities.
-Recently, Mikkola et al [51] proposed a method that uses pairwise comparisons or ranking techniques to learn a multivariate prior density, which could serve as an interesting avenue for future research.
 Other important design choices within our framework that require appropriate default settings include the specification of the discrepancy measure and the selection of the optimization approach, among other aspects.
-So far, in all simulation studies, we have used the maximum mean discrepancy [MMD, 36] to evaluate the discrepancy between the model-implied and expert-elicited statistics.
+So far, in all simulation studies, we have used the maximum mean discrepancy [MMD, 38] to evaluate the discrepancy between the model-implied and expert-elicited statistics.
 The only exception is the correlation information, for which we employed an L2 loss.
@@
 Comparing these approaches in terms of performance and the level of hyperparameter tuning required could offer valuable insights into the conditions under which optimization method might be more advantageous than the other.
+Finally, while our current methodology learns prior distributions based on expert data, it does currently not account for additional ‘meta’ uncertainty that arises during the elicitation process.
+This includes uncertainties introduced by the elicitation technique itself or the expert’s uncertainty regarding the ‘true’ quantity [see e.g., 24, 25, 71].
+We acknowledge that these are important sources of information that should be incorporated in future developments of our method.
+One promising approach is to view this problem from a Bayesian perspective and treat the proposed workflow as a Bayesian inference problem.
+This idea was already proposed by Mikkola et al [52] and referred to as the supra-Bayesian approach.
+We believe this is an important and valuable direction for the next steps in our elicitation method.
 Declarations
@@
 We thank Dmytro Perepolkin and Luna Fazio for their valuable comments and suggestions that greatly improved the manuscript.
-All authors gratefully acknowledge the computing time provided on the Linux HPC cluster at Technical University Dortmund (LiDO3), partially funded in the course of the Large-Scale Equipment Initiative by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as project 271512359.
 Funding
@@
 Data availability
-The simulation results are available at https://osf.io/xrzh6/.
+The simulation results are available on OSF at https://osf.io/xrzh6/.
 Materials availability
-Not applicable.
+Supplementary material is available on GitHub at https://github.com/florence-bockting/non-parametric-prior.
 Code availability
-[…]
+Code for running the simulations and plotting the results is available on GitHub at https://github.com/florence-bockting/non-parametric-prior.
+Simulations and most plots are based on/provided by our Python package elicito [5, v0.3.1].
+To ensure the reproducibility of our results, the specific version of elicito used in this study has been archived on Zenodo and is accessible via the following DOI https://doi.org/10.5281/zenodo.15241853.
 Author contribution
-Conceptualization: PB, SR, FB, Methodology: PB, FB, Software: FB, Writing (Original Draft): FB, Writing (Review & Editing): SR, PB, Visualization: FB, Supervision: PB, Funding acquisition: PB, SR
+Conceptualization: PB, SR, FB, Methodology: PB, FB, Software: FB, Writing (Original Draft): FB, Writing (Editing): FB, SR, PB, Writing (Review): FB, PB, Visualization: FB, Supervision: PB, Funding acquisition: PB
 Appendix A Simulation Studies
 A.1 Training time
-The average training time across the 30 replications was
-  • Binomial model: 24 minutes and 10 seconds ($\pm$ 33 seconds).
-  • Independent normal scenario: 66 minutes and 50 seconds ($\pm$ 5 min: 53 secs).
-  • Skewed normal scenario: 74 minutes and 34 seconds ($\pm$ 9 min: 16 secs).
-  • Correlated normal scenario: 73 minutes and 18 seconds ($\pm$ 1 min: 59 secs).
+The median training time across the 30 replications ($\pm$ sd of training time):
+  • Binomial model: 6.87 minutes ($\pm$ 1.45).
+  • Independent normal scenario: 10.18 minutes ($\pm$ 1.68).
+  • Skewed normal scenario: 14.45 minutes ($\pm$ 2.80).
+  • Correlated normal scenario: 10.10 minutes ($\pm$ 0.17).
 A.2 Independent normal scenario
@@
 In each row, the corresponding hyperparameter, is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line.
-The horizontal gray line is included as a reference to distinguish between the different model parameters.
-Figure 15: Normal model / Scenario 1 (M2) — Preliminary Convergence Check for all 30 replications: The upper plot displays the absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis.
+Figure 16: Normal model / Scenario 1 (M2) — Preliminary Convergence Check: Upper plot displays the absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis.
 The replications (i.e., seeds) are arranged in ascending order based on the magnitude of the absolute slope.
-The best-performing model, with an almost zero slope, is seed 13, while the worst-performing models, indicated by negative slopes, are seeds 11, 2, 12, and 28.
-The lower plot illustrates the loss trajectories of the best- and worst-performing seeds over the last 200 epochs.
-The loss updates for the final 100 epochs are highlighted in green for the best model and in blue for the worst models, with a red line segment indicating the respective slope.
-Figure 16: Normal model / Scenario 1 (M2) — Visual convergence check for seed=28: (a) Updating trajectory across epochs for various quantities of interest: the total loss (upper left), individual loss components (upper right), and the mean and standard deviation of the marginal priors (lower left and lower right, respectively).
-(b) Learned joint prior with marginal distributions along the diagonal and correlation information in the off-diagonal.
+The best-performing model is seed 14, while the worst-performing models are seeds 6, 26, 10, and 9.
+The lower plot illustrates the loss trajectories of the best- and worst-performing seeds over the last 300 epochs.
+The loss updates for the final 100 epochs are highlighted in green (best model) and blue (worst models), with a red line segment indicating the respective slope.
+Figure 17: Normal model / Scenario 1 (M2) — Visual convergence check 1: Updating trajectory across epochs for the total loss (left) and the individual (weighted) loss components (right).
+Figure 18: Normal model / Scenario 1 (M2) — Visual convergence check 2: Updating trajectory across epochs for the mean (left) and standard deviation (right) of the final learned marginal priors for all 30 replications.
+Figure 19: Normal model / Scenario 1 (M2) — Joint prior: Each replication is visualized in a different color.
 A.3 Skewed normal scenario
@@
 In each row, the corresponding hyperparameter, is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line.
-The horizontal gray line is included as a reference to distinguish between the different model parameters.
 Figure 21: Normal model / Scenario 2 (M3) — Preliminary Convergence Check for all 30 replications: The upper plot displays the absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis.
 The replications (i.e., seeds) are arranged in ascending order based on the magnitude of the absolute slope.
-The best-performing model, with an almost zero slope, is seed 21, while the worst-performing models, indicated by negative slopes, are seeds 15, 7, 8, and 27.
-The lower plot illustrates the loss trajectories of the best- and worst-performing seeds over the last 200 epochs.
-The loss updates for the final 100 epochs are highlighted in green for the best model and in blue for the worst models, with a red line segment indicating the respective slope.
-Figure 19: Normal model / Scenario 2 (M3) — Visual convergence check for seed=27: (a) Updating trajectory across epochs for various quantities of interest: the total loss (upper left), individual loss components (upper right), and the mean and standard deviation of the marginal priors (lower left and lower right, respectively).
-(b) Learned joint prior with marginal distributions along the diagonal and correlation information in the off-diagonal.
+The best-performing model is seed 13, while the four worst-performing models are seeds 4, 10, 3, and 16.
+The lower plot illustrates the loss trajectories of the best- and worst-performing seeds over the last 300 epochs.
+The loss updates for the final 100 epochs are highlighted in green (best seed) or blue (worst seeds).
+The red line segment indicates the respective slope.
+Figure 22: Normal model / Scenario 2 (M3) — Visual convergence check 1: Updating trajectory across epochs for the total loss (left) and the individual (weighted) loss components (right).
+Figure 23: Normal model / Scenario 2 (M3) — Visual convergence check 2: Updating trajectory across epochs for the mean (left) and standard deviation (right) of the marginal priors.
+Figure 24: Normal model / Scenario 2 (M3) — Visual convergence check 2: Joint prior distribution. Each replication is visualized in a different color.
 A.4 Correlated normal scenario
@@
 In each row, the corresponding hyperparameter, is varied across the range shown on the x-axis, while all other hyperparameters are held constant at their true values, indicated by the red vertical line.
-The horizontal gray line is included as a reference to distinguish between the different model parameters.
 Figure 26: Normal model / Scenario 3 (M4) — Preliminary Convergence Check for all 30 replications: The upper plot displays the absolute slope of the total loss trajectory (scaled by a factor of 100) over the last 100 epochs on the y-axis, with each replication shown along the x-axis.
 The replications (i.e., seeds) are arranged in ascending order based on the magnitude of the absolute slope.
-The best-performing model, with an almost zero slope, is seed 19, while the worst-performing models, indicated by negative slopes, are seeds 14, 5, 27, and 6.
+The best-performing model is seed 6, while the worst-performing models are seeds 11, 0, 13, and 20.
 The lower plot illustrates the loss trajectories of the best- and worst-performing seeds over the last 200 epochs.
 The loss updates for the final 100 epochs are highlighted in green for the best model and in blue for the worst models, with a red line segment indicating the respective slope.
-Figure 22: Normal model / Scenario 3 (M4) — Visual convergence check for seed=26: (a) Updating trajectory across epochs for various quantities of interest: the total loss (upper left), individual loss components (upper right), and the mean and standard deviation of the marginal priors (lower left and lower right, respectively).
-(b) Learned joint prior with marginal distributions along the diagonal and correlation information in the off-diagonal.
+Figure 27: Normal model / Scenario 3 (M4) — Visual convergence check 1: Updating trajectory across epochs for the total loss (left) and the individual (weighted) loss components (right).
+Figure 28: Normal model / Scenario 3 (M4) — Visual convergence check 1: Updating trajectory across epochs for the mean (left) and standard deviation (right) of the marginal priors.
+Figure 29: Normal model / Scenario 3 (M4) — Visual convergence check 2: Joint prior distribution. Each replication is visualized in a different color.